### v1

# MedChain: Bridging the Gap Between LLM Agents and Clinical Practice through Interactive Sequential Benchmarking

Abstract
Clinical decision making (CDM) is a complex, dynamic process crucial to healthcare delivery, yet it remains a significant challenge for artificial intelligence systems. While Large Language Model (LLM)-based agents have been tested on general medical knowledge using licensing exams and knowledge question-answering tasks, their performance in the CDM in real-world scenarios is limited due to the lack of comprehensive testing datasets that mirror actual medical practice. To address this gap, we present MedChain, a dataset of 12,163 clinical cases that covers five key stages of clinical workflow. MedChain distinguishes itself from existing benchmarks with three key features of real-world clinical practice: personalization, interactivity, and sequentiality. Further, to tackle real-world CDM challenges, we also propose MedChain-Agent, an AI system that integrates a feedback mechanism and a MCase-RAG module to learn from previous cases and adapt its responses. MedChain-Agent demonstrates remarkable adaptability in gathering information dynamically and handling sequential clinical tasks, significantly outperforming existing approaches. The relevant dataset and code will be released upon acceptance of this paper.
MedChain: Bridging the Gap Between LLM Agents and Clinical Practice through Interactive Sequential Benchmarking
Jie Liu1thanks: J. Liu ($\spadesuit$ $\clubsuit$ $\diamondsuit$), W. Wang ($\heartsuit$ $\clubsuit$) make equal contribution. $\spadesuit$: Conceptualization; $\heartsuit$: Implementation; $\clubsuit$: Writing; $\diamondsuit$: Visualization., Wenxuan Wang2∗, Zizhan Ma2, Guolin Huang3, SU Yihang2, Kao-Jung Chang4,5, Wenting Chen1, Haoliang Li1, Linlin Shen3, Michael Lyu2 1The City University of Hong Kong 2The Chinese University of Hong Kong 3Shenzhen University 4National Yang Ming Chiao Tung University 5Taipei Veterans General Hospital Project: https://github.com/ljwztc/MedChain
1 Introduction
At the intersection of artificial intelligence and healthcare lies one of medicine’s most complex challenges: Clinical Decision Making (CDM). In healthcare delivery, CDM demands not only the synthesis of diverse data sources and continuous assessment of evolving clinical scenarios, but also evidence-based judgments for diagnosis and treatment Sutton et al. (2020). While crucial for optimal patient care, this intricate process imposes significant cognitive demands on healthcare professionals, making it an ideal candidate for AI assistance Sendak et al. (2020).
Recent advances in Large Language Model (LLM)-based agents OpenAI (2023); Team et al. (2023); Gu et al. (2023); Shinn et al. (2024); Guan et al. (2023); Zhuang et al. have emerged as an effective solution for complex decision-making tasks, from software development Qian et al. (2024) to office automation Wang et al. (2024c). In the medical domain, these agents have demonstrated impressive performance on medical licensing exams Singhal et al. (2023); Pal et al. (2022) and knowledge-based assessments Gilson et al. (2023); Eriksen et al. (2023). While LLMs have consistently scored well above passing thresholds in these assessments Singhal et al. (2023), it is crucial to recognize that these assessments fall short of capturing the complexity of real-world CDM in three critical aspects.
Firstly, these benchmarks rarely account for patient-specific information such as past medical history and present illness Pal et al. (2022), which significantly influence clinical decisions in real clinical scenarios. This omission fails to capture the nuanced context that often shapes personalized diagnosis. Secondly, unlike real clinical scenarios where decisions build upon previous steps, existing benchmarks present clinical tasks as independent problems Schmidgall et al. (2024), missing the critical interdependencies in the diagnostic process. In reality, clinical decision-making is a sequential process where each step is contingent upon the preceding ones, and an error in one stage can profoundly impact subsequent decisions. Thirdly, most benchmarks present all relevant information upfront, providing a static, comprehensive dataset Tu et al. (2024). However, real clinical workflow demand multiple rounds of dynamic information gathering through ongoing patient interaction.
Figure 1: Demonstration of error propagation of CDM in MedChain. Starting with 2,362 initial cases, the diagram illustrates how diagnostic errors cascade through five clinical stages. Cases with incorrect diagnoses carry forward problematic information to subsequent stages, leading to a cumulative decrease in accuracy. After completing the treatment phase, we assess the correctness of decisions across all phases. Our MedChain-Agent achieves best performance in CDM comparing with other SOTA methods.
MedChain: To address these critical gaps, we introduce MedChain, a novel benchmark designed to evaluate LLM-based agents in real-world clinical scenarios. Specifically, MedChain comprises 12,163 diverse cases spanning 19 medical specialties and 156 sub-categories, including 7,338 medical images with corresponding reports. Each case progresses through five crucial stages: specialty referral, history-taking, examination, diagnosis, and treatment. Unlike existing benchmarks, MedChain uniquely emphasizes three key features. 1) Personalization: Each case incorporates detailed patient-specific information. At first, agents are provided with only the patient’s chief complaint and basic information. 2) Interactivity: Information must be actively gathered through dynamic consultation from patient. 3) Sequentiality: Decisions at each stage influence subsequent steps. Only after agent independently completes all five stages, the overall diagnostic process is evaluated.
MedChain-Agent: Given the novel features and challenges presented by this benchmark, existing agent frameworks struggle to address these issues adequately. To overcome these limitations, we propose MedChain-Agent, a multi-agent collaborative framework that enables LLMs with feedback mechanism and MedCase-RAG to dynamically gather information and handle sequential clinical tasks. Specifically, MedChain-Agent facilitates a synergistic interplay among three specialized agent types: General Agents for task-specific expertise, a Summarizing Agent for insight synthesis, and a Feedback Agent for iterative refinement. This multi-layered, iterative approach ensures decisions are products of thorough analysis and diverse perspectives. Additionally, to address the multifaceted nature of CDM, which demands the integration of evidence-based research, and patient-specific factors, we incorporate a novel MedCase-RAG module into our MedChain-Agent framework. Unlike conventional medical RAG methods, MedCase-RAG dynamically expand its database and employs a structured approach to data representation, mapping each medical case into a 12-dimensional feature vector. This system enables efficient retrieval of relevant cases and helps the model make informed decisions.
Our contributions are summarized as follows:
  • We represent the first effort to propose a CDM benchmark, MedChain, providing a holistic assessment of diagnostic capabilities of LLLM-based agents, closely reflecting real-world patient care.
  • We propose a multi-agent framework based on the characteristics of CDM. This system enables efficient retrieval of relevant cases and helps the model make informed decisions.
  • Through extensive experiments, we demonstrate the effectiveness of MedChain and the MedChain-Agent framework in improving clinical decision-making accuracy and reliability.
Table 1: Comparison of Benchmarks for LLM-based Agent. The MedChaincomposed of 12, 163 case with five sequential stages of the clinical workflow. It provide interaction environment for LLM-based agent evaluation.
Benchmark #Num Modality Personal Information Sequential Evaluation Interaction
MedQA 2018 1,273 Text
PubMedQA 2019 500 Text
MedMCQA 2022 193,155 Text
MMedBench 2024 53,566 Text
MedBench 2024b 300,901 Text
Asclepius 2024b 3,232 Text, Imaging
GMAI-MMBench 2024b 26,000 Text, Imaging
MedChain 12,163 Text, Imaging
2 Related Works
2.1 Evaluation of LLM in Medcine
Benchmarking plays a vital role as a key performance indicator, directing model improvements, pinpointing weaknesses, and shaping the course of model evolution. The evaluation of LLMs in medicine has primarily focused on testing general medical knowledge through structured assessments. Leading benchmarks such as MultiMedQA Singhal et al. (2023) integrate various medical QA datasets (e.g., MedQA Jin et al. (2021), MedMCQA Pal et al. (2022)), emphasizing performance on medical licensing examination materials. Other benchmarks like PubMedQA Jin et al. (2019) focus on research-oriented queries, while several Chinese medical benchmarks Wang et al. (2023); Cai et al. (2024) evaluate models through multiple-choice questions from medical licensing exams.
While these benchmarks effectively assess general medical knowledge, they fail to capture three critical aspects of real-world clinical decision-making (see Table 1), i.e., personalization in patient care, the interactive nature of clinical consultations, and the sequential dependency of medical decisions, where each step builds upon previous findings.
2.2 LLM-based Agent in Medicine
LLM-based agents have demonstrated significant potential across various medical applications, encompassing tasks such as medical examination questions, clinical diagnoses, and treatment plans. Recent research has explored different approaches: Agent Hospital Li et al. (2024) provides medical scenario simulation, while several frameworks Yue et al. (2024); Tang et al. (2024); Kim et al. (2024b) focus on specific medical stages with multi-agent architectures. Some works target specialized aspects, such as CoD Chen et al. (2024a) for interpretable diagnostics and Ehragent Shi et al. (2024) for electronic health records (EHRs) analysis. Others, like Almanac Copilot Zakka et al. (2024), assist clinicians with EMR-specific tasks. AI Hospital Fan et al. (2024) explores interactive clinical scenarios, but it falls short in handling multi-modal medical imaging and lacks a comprehensive benchmark for evaluating multi-agent performance. To enhance these agents’ capabilities, researchers have integrated Retrieval-Augmented Generation (RAG), as demonstrated by MIRAGE Xiong et al. (2024)’s search-enhanced framework and Medical Graph RAG Wu et al. (2024)’s knowledge-based approach.
However, current approaches face two major limitations. First, existing frameworks focus on isolated medical tasks rather than providing seamless integration across different clinical stages, making them insufficient for complex scenarios requiring effective inter-stage communication Gronauer and Diepold (2022). Second, current medical RAG systems’ reliance on chunk-based indexing leads to context inconsistency and computational inefficiencies Huang and Huang (2024); Gao et al. (2024), highlighting the need for more sophisticated approaches to medical knowledge integration.
3 MedChain Benchmark
Overview. We introduce MedChain, a comprehensive clinical decision-making benchmark designed to simulate real-world scenarios. Built upon 12,163 Electronic Health Records (EHRs) spanning 19 medical specialties and 156 sub-categories, including 7,338 medical images with reports, MedChain uniquely emphasizes three key characteristics:
  • Personalization: Each case includes detailed patient profiles that influence decision-making
  • Sequentiality: Cases involve multiple interconnected decision-making stages
  • Interactivity: Information must be actively gathered through dynamic consultation
Figure 2: MedChain Pipeline. The MedChain is composed of a sequential medical process, including specialty referral, history-taking , examination, diagnosis, and treatment.
3.1 Data Collection
Data Source and Processing. Our dataset is sourced from the Chinese medical website “iiYi”  [^1] https://www.iiyi.com, which provides over 20,000 validated clinical cases spanning 28 disease categories. These cases are verified by professional doctors and have undergone de-identification to ensure patient privacy. We obtained formal permission from the website administrators to use the data for scientific research purposes. Each case typically contains the patient’s chief complaint, medical history, examination results, treatment process, and other relevant information, which insure the personalization of MedChain. Following the government standards  [^2] National Standards for Medical Items and Medical Subject Headings  [^3] https://www.ncbi.nlm.nih.gov/mesh/1000048, we extracted and organized key information including patient basics, chief complaints, specialty referrals, examinations, imaging reports, diagnoses, and treatments. Cases with incomplete information were removed, resulting in 12,163 high-quality cases.
Quality Control. To ensure the highest standards of data integrity and clinical relevance in our benchmark, we implemented a rigorous quality control process involving a panel of five senior physicians, each with over 10 years of clinical experience. Our evaluation process examined a random sample of 6,000 cases (49.3% of the dataset). Each case was assessed on multiple dimensions including disease prevalence, clinical relevance, accuracy of patient history, appropriateness of diagnostic procedures, correctness of diagnosis, and suitability of treatment recommendations. We employed a standardized scoring system to quantify case quality across these dimensions and calculated inter-rater reliability using Cohen’s kappa coefficient. The quality assessment yielded impressive results, with an overall quality rate of 94.7% of evaluated cases meeting or exceeding our stringent quality thresholds. Dimension-specific quality rates ranged from 92.9% to 97.2%, demonstrating consistently high standards across all evaluation criteria. Our inter-rater reliability analysis produced a Cohen’s kappa coefficient of 0.82, indicating strong agreement among our expert reviewers. Cases that did not meet the thresholds (5.3%) underwent revision or were excluded from the final dataset to maintain benchmark integrity.
3.2 Clinical Workflow Simulation
Sequential Stages. MedChain simulates the complete clinical workflow, comprising five sequential tasks, each representing a different stage of the clinical decision-making process, as shown in Figure 2. The results from each stage serve as inputs for the subsequent stage, creating a dependency where later decisions are influenced by the quality of earlier ones. This design guarantees the sequentiality of MedChain, mimicking the interconnected nature of real-world clinical decision-making processes. The pipeline consists of: 1) Specialty Referral: Assessment of case urgency and appropriate department selection; 2) History-taking: Dynamic information gathering through doctor-patient dialogue; 3) Examination: Medical image analysis and report generation; 4) Diagnosis: Comprehensive diagnosis based on accumulated information; 5) Treatment: Treatment plan formulation considering patient-specific factors.
The constrcution pipeline of MedChain  is sketched out here. The details of benchmark can be refer to Appendix A. We demonstrate several examples in Appendix Figure 4.
Figure 3: MedChain-Agent framework. Depicts a cyclical feedback medical multi-task system, where decisions are supported by retrieving similar past cases from a medical database.
3.3 Interaction Environment
To simulate the doctor-patient consultation process in real medical scenarios, we have developed an interactive environment where the LLM-based agent under evaluation must actively gather information through dynamic interactions. We employ a local large language model, i.e., gemma2 (9b) Team et al. (2024), to serve as the patient agent. This patient agent is initialized with pre-defined case information but is unaware of the actual diagnosis, allowing it to provide symptom information and respond to inquiries in a manner that closely mimics real patient experiences. This setup ensures the interactivity for the MedChain benchmark, enabling it to evaluate the LLM’s ability to conduct effective medical consultations, ask pertinent questions, and gather crucial information for accurate diagnosis and treatment planning. The inspiration for this patient agent setup comes from standardized patients in medical education and clinical practice. Standardized patients are trained individuals who simulate real patients, allowing healthcare professionals to practice their clinical and communication skills in a controlled environment without the risks associated with real-patient interactions Barrows (1993).
3.4 Benchmark Evaluation
Given the complexity of medical decision-making, a simple binary evaluation of correct or incorrect is inadequate. To quantify the quality of CDM results, we’ve developed an evaluation system for each stage. The standard answer for each stage comprises multiple key points. The evaluated agent’s score increases with the number of standard points addressed, while irrelevant content leads to score reduction. To implement this approach, we employ the Intersection over Union (IoU) metric. For instance, in the treatment stage, the agent’s score improves as it proposes more reasonable treatment measures.
4 MedChain-Agent Framework
4.1 Multi-agent System with Feedback
The MedChain-Agent framework introduces a novel multi-agent system designed to simulate the complex, interconnected nature of medical decision-making processes. This system leverages the collective intelligence of multiple specialized agents, each contributing unique expertise to the overall decision-making process.
4.1.1 Agent Role
Our framework incorporates three distinct agent types, each fulfilling crucial roles in the medical decision-making pipeline:
General Agents: These specialized agents are recruited based on the specific requirements of each task. For instance, in Task 1 Specialty Referral, General Agents possess comprehensive knowledge of medical specialties and triage protocols, enabling them to effectively assess and route patients to appropriate departments. These agents engage in group discussions, simulating the collaborative nature of medical consultations.
Summarizing Agent: In each stage of the process, a summarizing agent consolidates the discussions and analyses of the general agents. This agent synthesizes the collective expertise into a coherent summary and delivers the final decision for each task, mimicking the role of a senior physician or department head within a medical context.
Feedback Agent: The feedback agent plays a crucial role in maintaining the quality and accuracy of decisions throughout the process. This agent evaluates the output of each task, providing constructive feedback and suggesting iterative improvements when necessary. By doing so, it creates a dynamic, self-correcting system that continually refines its decision-making capabilities.
4.1.2 Decision Making
With three types of agents, MedChain-Agent conducts a sophisticated decision-making process that mirrors real-world medical practice. General agents analyze patient information and engage in collaborative discussions, sharing assessments and recommendations. The summarizing agent then consolidates these insights, weighing different perspectives to formulate a preliminary decision. The feedback agent reviews this decision, evaluating its appropriateness and implications for subsequent tasks. If issues are identified, the feedback agent provides critiques and suggestions, initiating an iterative refinement process. General agents reconsider their assessments based on this feedback, while the summarizing agent refines the decision. This cycle continues until consensus is reached or a predefined number of iterations is completed, at which point the final decision is issued.
Table 2: Evaluation of various LLM-based agent in MedChain. The best performance for each task is highlighted in bold.
Framework Methods Specialty Referral History-taking Examination Diagnosis Treatment Average
Level 1 Level 2
Single-agent GPT-4o-mini 2023 0.5449 0.2871 0.3399 0.5112 0.4406 0.3930 0.4195
Claude-3.5-sonnet 2023 0.5681 0.3050 0.3562 0.5018 0.4599 0.4053 0.4327
LLaVA 2024a 0.3240 0.0730 0.3182 0.5165 0.4024 0.1060 0.2901
Qwen2 2024a 0.4975 0.2215 0.4226 0.4829 0.4818 0.2193 0.3876
InternVL2 2024a 0.4811 0.1935 0.4645 0.4490 0.4738 0.2903 0.3920
Multi-agent MedAgent 2023 0.3830 0.2039 0.4454 0.4456 0.4266 0.3673 0.3786
MDAgent 2024a 0.2398 0.1343 0.4240 0.4983 0.4289 0.3620 0.3479
MedChain-Agent 0.5873 0.3505 0.5836 0.6566 0.5218 0.4613 0.5269
4.2 MedCase-RAG
To enhance the decision-making capabilities of our multi-agent framework, we introduce MedCase-RAG, a novel Retrieval-Augmented Generation technique tailored for medical contexts.
Unlike conventional medical RAG methods that rely on unstructured medical QA databases, MedCase-RAG employs a structured approach to data representation. We map each medical case into a 12-dimensional feature vector, encompassing crucial attributes such as Age, Sex, Patient Description, Symptom Description, Patient History, and Patient Image. This structured representation allows for more comprehensive and standardized case encoding, facilitating more accurate and nuanced retrieval. Within this feature set, we identify "Symptom Description" as the most representative feature of a patient’s condition. This feature undergoes quantification through a Text Embedding model, serving as the primary key for dense retrieval tasks. Moreover, MedCase-RAG incorporates a dynamic database expansion mechanism to continually enrich its knowledge base. As patients complete their medical processes and recover, their case information is reintroduced into the database as pseudo-data. This approach allows the system to learn from new cases and adapt to evolving medical knowledge and practices.
When a new patient case is presented, MedCaseRAG performs a similarity search within the relevant medical department, using cosine similarity to identify the top three most similar cases, offering a more comprehensive basis for decision-making, allowing agents to consider various aspects and potential outcomes.
5 Experiments
5.1 Experimental Setup
We split the dataset into training, validation, and testing sets with a ratio of 7:1:2. For frameworks incorporating RAG techniques, both training and validation sets are used to construct the case retrieval database. Our study evaluates both single-agent and multi-agent systems. For single agents, we test two closed-source models (gpt-4o-mini OpenAI (2023), and claude-3.5-sonnet Team et al. (2023)) and three open-source models (InternVL2-8b Chen et al. (2023), llava-llama-3-8b-v1_1 Liu et al. (2024a), and Qwen2-7B-Instruct Wang et al. (2024a)), with model weights obtained from official Hugging Face repositories. In the multi-agent evaluation, we compare MedChain-Agent against MedAgent Tang et al. (2023) and MDAgent Kim et al. (2024a). All multi-agent framework are based on InternVL2-8b Chen et al. (2023). The deployment was conducted using the LMDeploy framework Contributors (2023). All tests executed on NVIDIA A100 GPUs featuring 80GB of memory. To enhance output stability and reliability across all experiments, we consistently set the temperature parameter to 0.
Table 3: Ablation Study for Key Components for MedChain-Agent. This table presents the performance impact of sequentially removing the Feedback mechanism and MedCase-RAG from the full MedChain-Agent framework.
Feedback MedCase-RAG Specialty referral History-taking Examination Diagnosis Treatment Average
Level 1 Level 2
0.5523 0.2228 0.3285 0.6369 0.4982 0.3915 0.4384
0.5739 0.2906 0.4222 0.6377 0.5086 0.4209 0.4757
0.5928 0.3353 0.5801 0.6488 0.5035 0.4568 0.5195
0.5873 0.3505 0.5836 0.6566 0.5218 0.4613 0.5269
5.2 Evaluation Metrics
To assess model performance, we employ diverse metrics tailored to each task. Accuracy is used for tasks with a single correct answer, such as Specialty Referral (Level 1). For tasks involving multiple key points, including Specialty Referral (Level 2), History-taking, and Treatment, we apply Intersection over Union (IoU) to capture the overlap between predicted and ground truth sets. Specifically for the Examination task, we utilize DocLens Xie et al. (2024) to evaluate the quality of image interpretation reports.
5.3 Experiment Result and Analysis
The results of our evaluation in the MedChain  are presented in Table 2. Our analysis yields three significant insights:
(1) Sequential decision-making tasks continue to pose significant challenges, even for advanced models. For instance, within the single-agent frameworks, GPT-4o-mini and InternVL2 achieve average scores of 0.4442 and 0.3920, respectively. These results indicate that despite their sophistication, these models struggle to maintain consistent performance across the sequential stages of clinical decision-making, highlighting the inherent difficulty of these tasks.
(2) Multi-agent frameworks based on InternVL2, such as MedAgent and MDAgent, exhibit inferior performance compared to their single-agent counterparts. This degradation suggests that traditional multi-agent approaches may exacerbate error propagation in sequential decision-making processes, leading to reduced overall performance. In contrast, our proposed MedChain-Agent significantly outperforms these multi-agent methods, achieving an average score of 0.5269. This improvement demonstrates that MedChain-Agent effectively mitigates error propagation, enhancing decision quality and reliability in clinical settings.
(3) The integration of the MedChain-Agent framework with open-source LLMs demonstrates significant superiority over proprietary models like GPT-4o-mini. The substantial performance gain observed with MedChain-Agent (average score of 0.5269) implies that our framework can leverage the strengths of open-source LLMs to achieve superior outcomes. This suggests that open-source models, when enhanced with our framework, are not only competitive but can also excel in handling intricate medical decision-making tasks.
5.4 Ablation Study
[…]
Setting Model Diagnosis Treatment
Full MedAgent 0.4266 0.3673
MDAgent 0.4289 0.3620
gpt-4o-mini 0.4406 0.3930
InternVL2 0.4759 0.4472
w/o Person. $\downarrow$ MedAgent 0.3274 $\downarrow$ 0.3754 $\uparrow$
[…]
w/o Inter. $\uparrow$ MedAgent 0.3315 $\downarrow$ 0.4109 $\uparrow$
MDAgent 0.4295 $\uparrow$ 0.3627 $\uparrow$
[…]
Three key characteristics for MedChain: To validate the effectiveness of personalization, interactivity, and sequentiality within our benchmark, we conduct ablation study as in Table 4. We systematically remove each characteristic and observe its impact on model performance across Diagnosis and Treatment tasks. 1) Removing patient-specific information (w/o Person.) consistently degrades performance across all models in diagnosis tasks (with drops ranging from 3.0% to 13.4%). This demonstrates that personalized information is crucial for accurate clinical decision-making. 2) Interestingly, when removing the sequential dependency between stages (w/o Seq.), most models show improved performance. This improvement actually validates the effectiveness of sequentiality in our benchmark, as it indicates that sequential decision-making poses greater challenges that better reflect real-world clinical scenarios. 3) Similarly, the improved performance observed after removing interactive information gathering (w/o Inter.) confirms the effectiveness of interactivity in our benchmark design. These results collectively suggest that while both sequentiality and interactivity make the benchmark more challenging, they are essential components that better simulate the complexity of real-world clinical decision-making processes.
Key components for MedChain-Agent: To assess the impact of individual components in our framework, we conducted ablation experiments, with results shown in Table 3. These findings highlight the crucial roles of both the Feedback mechanism and the MedCase-RAG module in enhancing overall performance. Each component independently contributes to significant improvements across various metrics. Notably, the MedCase-RAG module alone boosts Specialty Referral (Level 1) to 0.5928 and yields substantial gains in History-taking and Examination. When combined, Feedback and MedCase-RAG demonstrate a synergistic effect, achieving an impressive average score of 0.5269. This underscores the complementary nature of these components in optimizing our framework’s capabilities.
6 Conclusion
In this paper, we introduced MedChain, a novel benchmark for evaluating large language models in clinical decision-making, and MedChain-Agent, an innovative multi-agent framework designed to address the complexities of real-world medical scenarios. Our work bridges a critical gap between current AI capabilities and the nuanced realities of clinical practice by incorporating personalization, sequentiality, and interactivity into the evaluation process. MedChain-Agent’s sophisticated multi-agent system, enhanced by a feedback mechanism and the MedCase-RAG module, demonstrates unprecedented performance in sequential clinical decision-making tasks, even outperforming proprietary models when combined with open-source LLMs. By providing a more realistic and challenging evaluation framework, this research sets a new standard for evaluating and developing medical AI systems, paving the way for its responsible integration into clinical practice.
Acknowledgement
We are particularly indebted to the administrators of the iiyi website for their generosity in allowing us to utilize their data for our research purposes. We would like to acknowledge the assistance provided by Claude-3.5 in proofreading our manuscript for grammatical accuracy and in facilitating the creation of LaTeX tables.
Limitations
This paper has two primary limitations that offer avenues for future research:
  • Data Source Diversity: The MedChain benchmark is constructed from 12,163 electronic health records obtained from the Chinese medical website “iiYi." Although this dataset is extensive and covers 19 medical specialties and 156 sub-categories, it is derived from a single source. In our future work, we will incorporate additional data sources from different regions or healthcare systems to further enrich the dataset, providing a broader spectrum of clinical scenarios and enhancing the generalizability of the benchmark.
  • Patient Interaction Simulation: In our interactive environment, the patient responses during the history-taking stage are generated by the Gemma 2 language model. While this approach ensures consistency and control in evaluating the LLM-based agent, the real patient interactions can be more varied and complex. Future work could explore more advanced patient simulators or incorporate real dialogue data to capture a wider range of communication styles and behaviors.
Appendix for MedChain
Abstract.
Appendix A describes the process of standardizing and organizing the dataset for the MedChainbenchmark.
Appendix B provides a detailed explanation of the MedChain-Agent framework, including its implementation, feedback mechanism, and the novel Retrieval-Augmented Generation (RAG) approach used to enhance decision-making.
Appendix C lists the additional details for experiment.
Appendix A Benchmark Construction
A.1 Dataset Standardization
We employed a combination of large language models and human verification to label data across different tasks. Our methodology involves task-specific prompt construction and output matching to ensure data quality and diversity while maintaining alignment with real clinical case scenarios.
To ensure consistency and comparability across the benchmark, we standardized the classification of examination items into two main categories: Physical Examinations and Auxiliary Examinations. Physical Examinations include evaluations of various body systems and general health indicators, while Auxiliary Examinations encompass different imaging techniques and laboratory tests. We utilized GPT-4o to extract and classify examination items from each case, followed by manual verification to ensure accuracy. For medical imaging, we classified images into seven types, and manual review ensured the correctness of the classifications. Additionally, treatment items were extracted and categorized from each case. This standardization process ensures that the dataset is consistent, facilitating accurate and comparable evaluations of LLM performance. Figure 5 shows the statistics of different departments and Figure 4 demonstrate a case after standardization.
Figure 4: Case presentation "77_Ovarian Carcinoid with Mature Cystic Teratoma: A Case Report."
Figure 5: Multi-departmental Distribution
A.2 Tasks Details
This section provides a comprehensive description of the five specific tasks that make up the MedChainbenchmark. Each subsection elaborates on a particular task, detailing its input, output, and evaluation methods. Figure 6 visualizes the differences between MedChainand other methods.
A.2.1 Task 1: Specialty referral
The specialty referral task evaluates the LLM’s ability to assess the urgency of a patient’s condition and determine the appropriate department based on the patient’s chief complaint. The input consists of the patient’s chief complaint, and the output space includes 19 first-level departments and 156 second-level departments. The LLM must first assign the patient to one of the 19 first-level departments, then to one or more of the 156 second-level departments based on the primary symptoms. Evaluation metrics include accuracy for first-level department assignment and Intersection over Union (IoU) for second-level department assignment. IoU is used for second-level departments to account for the possibility of multiple correct assignments and to reflect partial correctness, which can occur in complex cases. This metric better captures the nuanced nature of departmental referrals in clinical practice.
A.2.2 Task 2: History-taking
The history-taking task is designed to simulate doctor-patient communication, where the goal is to obtain relevant information and infer necessary examination items. We employ a multi-agent system to evaluate this process:
  • Doctor Agent: The LLM being evaluated plays this role, asking questions and suggesting examinations based on the patient’s responses.
  • Patient Agent: A local large model (such as Google’s gemma2/9b) simulates the patient, responding based on pre-defined case information. This includes the patient’s chief complaint, medical history, and examination results.
  • Extraction Agent: This agent extracts examination items that the Doctor Agent inquires about or suggests from the dialogue between the Doctor and Patient Agents. These extracted items form the output for comparison with the ground truth examination set.
The evaluation metric is the IoU between the examination items extracted by the Extraction Agent (based on the Doctor Agent’s inquiries and suggestions) and the ground truth set. This multi-agent design simulates a realistic clinical history-taking process, allowing the LLM to demonstrate its ability to ask relevant follow-up questions, interpret patient responses, and determine appropriate examinations. The use of a local large model as the Patient Agent ensures consistency in evaluations and improves the reproducibility of the benchmark.
A.2.3 Task 3: Examination
The examination task assesses the LLM’s ability to analyze medical images and generate corresponding image reports. The input consists of the medical images from each case, and the output is a free-text image report. This task tests the LLM’s capability to interpret visual medical data and articulate findings in a clear, professional manner. Given the complexity of medical reports, conventional NLP metrics such as BLEU Papineni et al. (2002) and BERTScore Zhang et al. (2019) are inadequate for evaluating this task, as they do not capture whether the generated report aligns with the ground truth in terms of medical claims. Instead, we adopt the Claim Recall metric proposed in DocLens Xie et al. (2024) to evaluate the completeness of the generated report. First, GPT-4o-mini is used to extract a list of claims from the original ground truth report. Then, GPT-4o-mini assesses whether the generated report entails these reference claims, and the recall score measures the proportion of claims that are correctly reflected in the generated report. The prompts used for this evaluation can be found in Figure 7.
A.2.4 Task 4: Diagnosis
The diagnosis task evaluates the LLM’s ability to identify the patient’s illness based on the provided symptoms and examination results. The input includes the patient’s chief complaint, medical history, examination results, and the department visited. The output is a diagnosis in free-text format. We use a five-level grading system to evaluate the accuracy of the diagnosis, ranging from completely inaccurate to completely accurate. This graded system allows for a more nuanced evaluation of the LLM’s diagnostic capabilities, as diagnoses can be partially correct or expressed in different ways. The system accounts for the complex nature of medical diagnoses, where multiple factors may contribute to the patient’s condition. The five-level grading system is listed as following:
  1. 1.
    Completely inaccurate (diagnosis completely inconsistent with the actual diagnosis)
  2. 2.
    Mostly inaccurate (most of the diagnosis inconsistent with the actual diagnosis)
  3. 3.
    Partially accurate (part of the diagnosis consistent with the actual diagnosis, but overall large discrepancy)
  4. 4.
    Mostly accurate (most of the diagnosis consistent with the actual diagnosis)
  5. 5.
    Completely accurate (diagnosis fully consistent with the actual diagnosis)
A.2.5 Task 5: Treatment
The treatment task involves formulating a treatment plan based on the patient’s diagnosis and examination results. The input includes the patient’s chief complaint, medical history, examination results, and diagnosis. The output is a set of proposed treatment items. The evaluation metric is the IoU between the proposed treatments and the ground truth treatment set. Since multiple treatment options may be appropriate, IoU allows for partial credit when the LLM suggests a subset of the recommended treatments or proposes additional reasonable treatments that are not part of the ground truth. This task tests the LLM’s ability to synthesize the gathered information and apply medical knowledge to develop an appropriate treatment plan, considering factors such as drug interactions, treatment protocols, and patient-specific considerations.
Figure 6: Comparison between our benchmark method and previous benchmarks. In Figure (a), the benchmark methods used in previous work are shown, which are based on specific domain problems and cannot achieve a complete end-to-end medical task workflow. In Figure (b), our proposed benchmark is presented, where we effectively accomplish the full workflow of medical tasks through our proposed MCase-RAG method and the ACFM mechanism.
Figure 7: The prompt for Extract Reference Claim and Entailment Evaluation.
A.3 Examination Items
Physical Examination: General examination (including height, weight, temperature, blood pressure, pulse, etc.), head, eyes, ears, nose and throat examination, neck examination (including thyroid, cervical lymph nodes), chest examination (including lungs, heart), abdominal examination, spine and limb examination, skin examination, neurological examination, urogenital system examination.
Auxiliary Examinations: X-ray, MRI, CT, ultrasound, nuclear medicine imaging, blood tests, urine tests, stool tests, endoscopy, pathological examination.
A.4 Treatment Items
Surgery, interventional therapy, medication, chemotherapy, antibiotic therapy, radiation therapy, physical therapy, immunotherapy, psychological therapy, traditional Chinese medicine, gene therapy.
Figure 8: The specific agent design for each task in our MedChain-Agent and the ACFM of the entire process are presented.
Appendix B MedChain-Agent
B.1 Implementation Details
The overall process and output flow of our framework are illustrated in Figure 8. The patient first describes their condition, after which Task 1 performs specialty referral, assigning the patient to the appropriate primary and secondary departments. Once Task 1 is completed, the doctor group consists of specialists from the corresponding primary department, enabling better analytical decision-making. Task 2 involves these specialists prescribing suitable medical examinations based on the patient’s current situation. Task 3 focuses on analyzing the patient’s imaging data to generate a detailed and comprehensive imaging report. Task 4 consists of the specialists conducting a comprehensive analysis based on the results from Task 2 and the imaging report from Task 3 to provide an initial diagnosis for the patient. Finally, Task 5 involves the specialists formulating an appropriate treatment plan based on all the previous tasks’ results and analyses.
Additionally, in the aforementioned description, the decision-making and analytical processes for all tasks are conducted through discussions within the Doctor Group. Specifically, three specialists from the same department first engage in discussions and analysis, after which they summarize their findings and results to the final medical generalist (primary care physician) for the final decision. At each step, relevant similar cases are retrieved from our medical database to assist in decision-making. For Task 1, we will divide the tasks of identifying primary and secondary departments into two subtasks based on the Chain of Thought (CoT) approach: first identifying the primary department, followed by identifying the secondary department. Similarly, for Task 2, we first prescribe specific tests before additional examinations, optimizing the decision-making process.
Once the patient completes the entire medical process, a comprehensive treatment follow-up record will be compiled, including treatment outcomes and reports from each stage, along with the patient’s information. This data will then be reintroduced into our medical database. Since these data have not been fully validated (we do not know how reasonable the decisions made for the current cases are), we will treat this data as pseudo-data based on the principles of semi-supervised learning and assign it a lower priority. This means that we will first retrieve data from completely validated sources, and if retrieval fails or data is insufficient, we will resort to retrieving from the pseudo-data.
B.2 Feedback Mechanism Details
To date, no work has proposed a multi-agent framework for simulating the entire medical process. Although existing frameworks demonstrate efficiency in specific medical tasks, their performance is suboptimal when linking various medical stages together. In sequential task scenarios, simply concatenating individual tasks is not feasible. Therefore, our full-process task can be regarded as a multi-sequence task, which necessitates a deeper exploration of the issue of error propagation. When a problem arises in the first task, subsequent tasks will analyze and make decisions based on erroneous results, potentially leading to severe impacts on the entire process.
To address this issue, we have introduced a feedback mechanism within the full-process framework, tightly connecting the current task with all subsequent tasks (as shown in Figure 6). After the current task is completed, the output results are evaluated by the physicians of the subsequent tasks (Feedback Judgement Group). Only when the results are confirmed to be accurate will the process advance to the next task; if issues are identified, the reasons for the errors and improvement suggestions will be output and fed back to the current task for re-discussion and decision-making. This process will continue iteratively until consensus is reached on the results or the maximum number of discussion rounds is achieved.
By implementing a feedback mechanism between each task, we can effectively manage the propagation of errors while significantly enhancing collaboration and communication among agent groups. This mechanism encourages agents to share information in real-time, improving mutual understanding and allowing for rapid strategic adjustments when issues are identified, thereby enhancing the overall flexibility and adaptability of the system. Each agent can better respond to changes in a dynamic environment, facilitating more efficient medical services. This feedback-driven collaborative model lays the foundation for the efficient operation of multi-agent systems, aiding in the provision of precise and reliable medical decision support in complex clinical environments.
B.3 RAG Details
We performed data restructuring on the original dataset, where each case is mapped into a 12-dimensional feature vector using attributes like Age, Sex, Patient Description, Symptom Description, Patient History, and Patient Image. Among these, "Symptom Description" is identified as the most representative feature of the current patient and is processed through a Text Embedding model for quantification, which is stored in the database for subsequent dense retrieval tasks.
Previous Medical-RAG methods relied on medical question-answering (QA) databases and predominantly used chunked indexing for retrieval. Our approach differs in several ways (as shown in Figure 5). Firstly, the content of our foundational medical database is distinct; while most methods have built their databases using medical QA data, ours utilizes a tree structure. We initially categorize patient information according to primary medical departments (in the experimental section, we categorized into 19 primary departments) and then extract and map patient information into a two-dimensional feature representation, as illustrated in Figure 4. This storage design allows for better retention of the patient’s crucial clinical information to assist in decision-making. Secondly, our retrieval method also diverges from theirs. In our retrieval process, we simulate the everyday practice of physicians. When faced with challenging cases that require decision-making, doctors often refer to past cases as references to enhance current decisions and judgments. We treat these two-dimensional features as the minimal unit of a case, using “Symptom description” as the basis for retrieval. When a new patient arrives, we extract and refine their “Symptom description” feature (approximately 70 characters). We employ direct quantitative matching without tokenizing this feature, using a Text Embedding model for quantification. The resulting feature vector is then compared to each case in our database corresponding to the relevant department through cosine similarity calculations, selecting the top three cases with the highest symptom similarity as the current retrieval results. These results, combined with the current case features, are forwarded to subsequent agents for comprehensive decision-making and judgment, enhancing the agents’ output. Since our retrieval method does not utilize document chunking but rather employs direct matching, it retains more matching information, leading to improved matching accuracy.
B.4 Task Prompts
[…]
Figure 9: The prompt for Task 1 in the MedChain-Agent .
Figure 10: The prompt for Task 2 in the MedChain-Agent .
Figure 11: The prompt for Task 3 in the MedChain-Agent .
Figure 12: The prompt for Task 4 and Task5 in the MedChain-Agent .
Appendix C Supplementary for Experiment
Tasks 1 and 2 are subdivided based on the Chain of Thought (CoT) reasoning approach for decision-making. Consequently, Task 1 is split into "specialty referral (Level 1)" and "specialty referral (Level 2)" to evaluate performance at both primary and secondary department levels. Task 2 is divided into "Physical" and "Ancillary," representing performance in physical examinations and ancillary tests. or both ablation and comparative experiments, we split our dataset into training and test sets in a 7:3 ratio. We extracted 12-dimensional features from each training set case and stored them in our Medical Dataset for subsequent retrieval tasks.

### v2

@@
-MedChain: Bridging the Gap Between LLM Agents and Clinical Practice through Interactive Sequential Benchmarking
+MedChain: Bridging the Gap Between LLM Agents and Clinical Practice with Interactive Sequence
 Abstract
 Clinical decision making (CDM) is a complex, dynamic process crucial to healthcare delivery, yet it remains a significant challenge for artificial intelligence systems.
-While Large Language Model (LLM)-based agents have been tested on general medical knowledge using licensing exams and knowledge question-answering tasks, their performance in the CDM in real-world scenarios is limited due to the lack of comprehensive testing datasets that mirror actual medical practice.
+While Large Language Model (LLM)-based agents have been tested on general medical knowledge using licensing exams and knowledge question-answering tasks, their performance in the CDM in real-world scenarios is limited due to the lack of comprehensive benchmark that mirror actual medical practice.
 To address this gap, we present MedChain, a dataset of 12,163 clinical cases that covers five key stages of clinical workflow.
 MedChain distinguishes itself from existing benchmarks with three key features of real-world clinical practice: personalization, interactivity, and sequentiality.
-Further, to tackle real-world CDM challenges, we also propose MedChain-Agent, an AI system that integrates a feedback mechanism and a MCase-RAG module to learn from previous cases and adapt its responses.
+Further, to tackle real-world CDM challenges, we also propose MedChain-Agent, an AI system that integrates a feedback mechanism and a MedCase-RAG module to learn from previous cases and adapt its responses.
 MedChain-Agent demonstrates remarkable adaptability in gathering information dynamically and handling sequential clinical tasks, significantly outperforming existing approaches.
 The relevant dataset and code will be released upon acceptance of this paper.
-MedChain: Bridging the Gap Between LLM Agents and Clinical Practice through Interactive Sequential Benchmarking
-Jie Liu1thanks: J.
-Liu ($\spadesuit$ $\clubsuit$ $\diamondsuit$), W.
-Wang ($\heartsuit$ $\clubsuit$) make equal contribution.
-$\spadesuit$: Conceptualization; $\heartsuit$: Implementation; $\clubsuit$: Writing; $\diamondsuit$: Visualization., Wenxuan Wang2∗, Zizhan Ma2, Guolin Huang3, SU Yihang2, Kao-Jung Chang4,5, Wenting Chen1, Haoliang Li1, Linlin Shen3, Michael Lyu2 1The City University of Hong Kong 2The Chinese University of Hong Kong 3Shenzhen University 4National Yang Ming Chiao Tung University 5Taipei Veterans General Hospital Project: https://github.com/ljwztc/MedChain
 1 Introduction
 At the intersection of artificial intelligence and healthcare lies one of medicine’s most complex challenges: Clinical Decision Making (CDM).
-In healthcare delivery, CDM demands not only the synthesis of diverse data sources and continuous assessment of evolving clinical scenarios, but also evidence-based judgments for diagnosis and treatment Sutton et al.
-(2020).
-While crucial for optimal patient care, this intricate process imposes significant cognitive demands on healthcare professionals, making it an ideal candidate for AI assistance Sendak et al.
-(2020).
-Recent advances in Large Language Model (LLM)-based agents OpenAI (2023); Team et al.
-(2023); Gu et al.
-(2023); Shinn et al.
-(2024); Guan et al.
-(2023); Zhuang et al.
-have emerged as an effective solution for complex decision-making tasks, from software development Qian et al.
-(2024) to office automation Wang et al.
-(2024c).
-In the medical domain, these agents have demonstrated impressive performance on medical licensing exams Singhal et al.
-(2023); Pal et al.
-(2022) and knowledge-based assessments Gilson et al.
-(2023); Eriksen et al.
-(2023).
-While LLMs have consistently scored well above passing thresholds in these assessments Singhal et al.
-(2023), it is crucial to recognize that these assessments fall short of capturing the complexity of real-world CDM in three critical aspects.
-Firstly, these benchmarks rarely account for patient-specific information such as past medical history and present illness Pal et al.
-(2022), which significantly influence clinical decisions in real clinical scenarios.
+In healthcare delivery, CDM demands not only the integration of diverse data sources and continuous assessment of evolving clinical scenarios, but also evidence-based judgments for diagnosis and treatment sutton2020overview .
+While crucial for optimal patient care, this intricate process imposes significant cognitive demands on healthcare professionals, making it an ideal candidate for AI assistance sendak2020human .
+Recent advances in Large Language Model (LLM)-based agents openai2023gpt4v ; team2023gemini ; gu2023don ; shinn2024reflexion ; guan2023leveraging ; zhuangtoolchain have emerged as an effective solution for complex decision-making tasks, from software development qian2024chatdev to office automation wang2024officebench .
+In the medical domain, these LLMs have demonstrated impressive performance on medical licensing exams singhal2023large ; pal2022medmcqa and knowledge-based assessments gilson2023does ; eriksen2023use ; liu2025comprehensive .
+While LLMs have consistently scored well above passing thresholds in these benchmark singhal2023large , it is crucial to recognize that these assessments fall short of capturing the complexity of real-world CDM, where errors can cascade through multiple decision stages, as illustrated in Figure 1.
+Based on our analysis, CDM exhibits three key characteristics.
+Firstly, these benchmarks rarely account for patient-specific information such as past medical history and present illness pal2022medmcqa , which significantly influence clinical decisions in real clinical scenarios.
 This omission fails to capture the nuanced context that often shapes personalized diagnosis.
-Secondly, unlike real clinical scenarios where decisions build upon previous steps, existing benchmarks present clinical tasks as independent problems Schmidgall et al.
-(2024), missing the critical interdependencies in the diagnostic process.
+Secondly, unlike real clinical scenarios where decisions build upon previous steps, existing benchmarks present clinical tasks as independent problems schmidgall2024agentclinic , missing the critical interdependencies in the diagnostic process.
 In reality, clinical decision-making is a sequential process where each step is contingent upon the preceding ones, and an error in one stage can profoundly impact subsequent decisions.
-Thirdly, most benchmarks present all relevant information upfront, providing a static, comprehensive dataset Tu et al.
-(2024).
+Thirdly, most benchmarks present all relevant information upfront, providing a static, and comprehensive dataset tu2024towards .
 However, real clinical workflow demand multiple rounds of dynamic information gathering through ongoing patient interaction.
@@
 Cases with incorrect diagnoses carry forward problematic information to subsequent stages, leading to a cumulative decrease in accuracy.
-After completing the treatment phase, we assess the correctness of decisions across all phases.
+After completing the treatment phase, we count cases that maintain correctness through each consecutive phase up to the each stage.
 Our MedChain-Agent achieves best performance in CDM comparing with other SOTA methods.
@@
 3) Sequentiality: Decisions at each stage influence subsequent steps.
-Only after agent independently completes all five stages, the overall diagnostic process is evaluated.
+Only after agent sequentially completes all five stages, the overall diagnostic process is evaluated.
 MedChain-Agent: Given the novel features and challenges presented by this benchmark, existing agent frameworks struggle to address these issues adequately.
@@
   • We represent the first effort to propose a CDM benchmark, MedChain, providing a holistic assessment of diagnostic capabilities of LLLM-based agents, closely reflecting real-world patient care.
-  • We propose a multi-agent framework based on the characteristics of CDM. This system enables efficient retrieval of relevant cases and helps the model make informed decisions.
-  • Through extensive experiments, we demonstrate the effectiveness of MedChain and the MedChain-Agent framework in improving clinical decision-making accuracy and reliability.
-Table 1: Comparison of Benchmarks for LLM-based Agent.
-The MedChaincomposed of 12, 163 case with five sequential stages of the clinical workflow.
-It provide interaction environment for LLM-based agent evaluation.
-Benchmark #Num Modality Personal Information Sequential Evaluation Interaction
-MedQA 2018 1,273 Text
-PubMedQA 2019 500 Text
-MedMCQA 2022 193,155 Text
-MMedBench 2024 53,566 Text
-MedBench 2024b 300,901 Text
-Asclepius 2024b 3,232 Text, Imaging
-GMAI-MMBench 2024b 26,000 Text, Imaging
-MedChain 12,163 Text, Imaging
+  • We propose a multi-agent framework based on the characteristics of CDM, called MedChain-Agent.
+This system enables efficient retrieval of relevant cases and helps the model make informed decisions.
+  • Through extensive experiments, we compare the performance of existing works on MedChain and the superiority of MedChain-Agent in CDM and realibility.
 2 Related Works
@@
 Benchmarking plays a vital role as a key performance indicator, directing model improvements, pinpointing weaknesses, and shaping the course of model evolution.
-The evaluation of LLMs in medicine has primarily focused on testing general medical knowledge through structured assessments.
-Leading benchmarks such as MultiMedQA Singhal et al.
-(2023) integrate various medical QA datasets (e.g., MedQA Jin et al.
-(2021), MedMCQA Pal et al.
-(2022)), emphasizing performance on medical licensing examination materials.
-Other benchmarks like PubMedQA Jin et al.
-(2019) focus on research-oriented queries, while several Chinese medical benchmarks Wang et al.
-(2023); Cai et al.
-(2024) evaluate models through multiple-choice questions from medical licensing exams.
-While these benchmarks effectively assess general medical knowledge, they fail to capture three critical aspects of real-world clinical decision-making (see Table 1), i.e., personalization in patient care, the interactive nature of clinical consultations, and the sequential dependency of medical decisions, where each step builds upon previous findings.
+The evaluation of LLMs in medicine has primarily focused on testing general medical knowledge through structured assessments ma2025beyond .
+Leading benchmarks such as MultiMedQA singhal2023large integrate various medical QA datasets (e.g., MedQA jin2021disease , MedMCQA pal2022medmcqa ), emphasizing performance on medical licensing examination materials.
+Other benchmarks like PubMedQA jin2019pubmedqa focus on research-oriented queries, while several Chinese medical benchmarks wang2023cmb ; cai2024medbench evaluate models through multiple-choice questions from medical licensing exams.
+While these benchmarks effectively assess general medical knowledge, they fail to capture three critical aspects of real-world clinical decision-making (see Appendix Table 6), i.e., personalization in patient care, the interactive nature of clinical consultations, and the sequential dependency of medical decisions, where each step builds upon previous findings.
+Recently, several benchmarks wu2025towards ; ouyang2024climedbench ; liu2025asclepius ; liang2024wsi ; li2025towards ; ding2025eagle have been proposed to evaluate LLMs across diverse clinical scenarios and tasks, including information extraction, text summarization, and clinical outcome prediction.
+However, these benchmarks primarily consist of independent question-answer pairs, where each task is evaluated in isolation.
+Sequential decision-making is critical in medical practice, as each patient experiences a continuous journey from initial triage through treatment to recovery.
+Therefore, it is essential to evaluate how LLMs perform throughout this entire clinical pipeline.
+Our work distinguishes itself from existing benchmarks by focusing on sequential decision-making within interactive environments, specifically evaluating LLM performance in realistic clinical scenarios that require executing the complete patient care workflow.
 2.2 LLM-based Agent in Medicine
-LLM-based agents have demonstrated significant potential across various medical applications, encompassing tasks such as medical examination questions, clinical diagnoses, and treatment plans.
-Recent research has explored different approaches: Agent Hospital Li et al.
-(2024) provides medical scenario simulation, while several frameworks Yue et al.
-(2024); Tang et al.
-(2024); Kim et al.
-(2024b) focus on specific medical stages with multi-agent architectures.
-Some works target specialized aspects, such as CoD Chen et al.
-(2024a) for interpretable diagnostics and Ehragent Shi et al.
-(2024) for electronic health records (EHRs) analysis.
-Others, like Almanac Copilot Zakka et al.
-(2024), assist clinicians with EMR-specific tasks.
-AI Hospital Fan et al.
-(2024) explores interactive clinical scenarios, but it falls short in handling multi-modal medical imaging and lacks a comprehensive benchmark for evaluating multi-agent performance.
-To enhance these agents’ capabilities, researchers have integrated Retrieval-Augmented Generation (RAG), as demonstrated by MIRAGE Xiong et al.
-(2024)’s search-enhanced framework and Medical Graph RAG Wu et al.
-(2024)’s knowledge-based approach.
+LLM-based agents have demonstrated significant potential across various medical applications wang-etal-2025-survey ; wsiagent , encompassing tasks such as medical examination questions, clinical diagnoses, and treatment plans.
+Recent research has explored different approaches: Agent Hospital agenthospitall provides medical scenario simulation, while several frameworks clinicalagent ; MEDAGENTS ; mdagent focus on specific medical stages with multi-agent architectures.
+Some works target specialized aspects, such as CoD chen2024cod for interpretable diagnostics and Ehragent shi2024ehragent for electronic health records (EHRs) analysis.
+Others, like Almanac Copilot zakka2024almanac , assist clinicians with EMR-specific tasks.
+AI Hospital fan2024ai explores interactive clinical scenarios, but it falls short in handling multi-modal medical imaging and lacks a comprehensive benchmark for evaluating multi-agent performance.
+To enhance these agents’ capabilities, researchers have integrated Retrieval-Augmented Generation (RAG), as demonstrated by MIRAGE MIRAGE ’s search-enhanced framework and Medical Graph RAG GraphRAG ’s knowledge-based approach.
 However, current approaches face two major limitations.
-First, existing frameworks focus on isolated medical tasks rather than providing seamless integration across different clinical stages, making them insufficient for complex scenarios requiring effective inter-stage communication Gronauer and Diepold (2022).
-Second, current medical RAG systems’ reliance on chunk-based indexing leads to context inconsistency and computational inefficiencies Huang and Huang (2024); Gao et al.
-(2024), highlighting the need for more sophisticated approaches to medical knowledge integration.
+First, existing frameworks focus on isolated medical tasks rather than providing seamless integration across different clinical stages, making them insufficient for complex scenarios requiring effective inter-stage communication coorapation .
+Second, current medical RAG systems’ reliance on chunk-based indexing leads to context inconsistency and computational inefficiencies RAGsurvey1 ; RAGsurvey2 , highlighting the need for more sophisticated approaches to medical knowledge integration.
+Figure 2: MedChain Pipeline. The MedChain is composed of a sequential medical process, including specialty referral, history-taking , examination, diagnosis, and treatment.
 3 MedChain Benchmark
@@
   • Interactivity: Information must be actively gathered through dynamic consultation
-Figure 2: MedChain Pipeline. The MedChain is composed of a sequential medical process, including specialty referral, history-taking , examination, diagnosis, and treatment.
 3.1 Data Collection
@@
 Our evaluation process examined a random sample of 6,000 cases (49.3% of the dataset).
-Each case was assessed on multiple dimensions including disease prevalence, clinical relevance, accuracy of patient history, appropriateness of diagnostic procedures, correctness of diagnosis, and suitability of treatment recommendations.
-We employed a standardized scoring system to quantify case quality across these dimensions and calculated inter-rater reliability using Cohen’s kappa coefficient.
-The quality assessment yielded impressive results, with an overall quality rate of 94.7% of evaluated cases meeting or exceeding our stringent quality thresholds.
+We developed a standardized scoring system that presents physicians with comprehensive case information alongside six binary quality dimensions: disease prevalence, clinical relevance, accuracy of patient history, appropriateness of diagnostic procedures, correctness of diagnosis, and suitability of treatment recommendations.
+Physicians evaluate each dimension through yes/no responses, with cases satisfying all dimensions considered valid.
+The quality assessment yielded strong results, with 94.7% of evaluated cases meeting or exceeding our quality thresholds.
 Dimension-specific quality rates ranged from 92.9% to 97.2%, demonstrating consistently high standards across all evaluation criteria.
-Our inter-rater reliability analysis produced a Cohen’s kappa coefficient of 0.82, indicating strong agreement among our expert reviewers.
-Cases that did not meet the thresholds (5.3%) underwent revision or were excluded from the final dataset to maintain benchmark integrity.
+Inter-rater reliability analysis produced a Cohen’s kappa coefficient of 0.82, indicating substantial agreement among our expert reviewers.
+Cases that failed to meet the thresholds (5.3%) were either revised or excluded from the final dataset to maintain benchmark integrity.
 3.2 Clinical Workflow Simulation
@@
 The pipeline consists of: 1) Specialty Referral: Assessment of case urgency and appropriate department selection; 2) History-taking: Dynamic information gathering through doctor-patient dialogue; 3) Examination: Medical image analysis and report generation; 4) Diagnosis: Comprehensive diagnosis based on accumulated information; 5) Treatment: Treatment plan formulation considering patient-specific factors.
-The constrcution pipeline of MedChain  is sketched out here. The details of benchmark can be refer to Appendix A. We demonstrate several examples in Appendix Figure 4.
+The construction pipeline of MedChain  is sketched out here.
+The standardization process and the format of each pipeline can be refer to Appendix A.
+We demonstrate an example in Appendix Figure 4 and Figure 5.
 Figure 3: MedChain-Agent framework. Depicts a cyclical feedback medical multi-task system, where decisions are supported by retrieving similar past cases from a medical database.
 3.3 Interaction Environment
-To simulate the doctor-patient consultation process in real medical scenarios, we have developed an interactive environment where the LLM-based agent under evaluation must actively gather information through dynamic interactions.
-We employ a local large language model, i.e., gemma2 (9b) Team et al.
-(2024), to serve as the patient agent.
-This patient agent is initialized with pre-defined case information but is unaware of the actual diagnosis, allowing it to provide symptom information and respond to inquiries in a manner that closely mimics real patient experiences.
-This setup ensures the interactivity for the MedChain benchmark, enabling it to evaluate the LLM’s ability to conduct effective medical consultations, ask pertinent questions, and gather crucial information for accurate diagnosis and treatment planning.
-The inspiration for this patient agent setup comes from standardized patients in medical education and clinical practice.
-Standardized patients are trained individuals who simulate real patients, allowing healthcare professionals to practice their clinical and communication skills in a controlled environment without the risks associated with real-patient interactions Barrows (1993).
+To simulate authentic doctor-patient consultations, we developed an interactive environment where LLM-based agents must actively gather information through dynamic interactions.
+We employ Gemma2 (9b) team2024gemma as the patient agent, initializing it with pre-defined case information while withholding the actual diagnosis.
+Each tested LLM serves as a doctor simulator.
+This setup enables the agent to provide symptom information and respond to inquiries in a manner that mirrors real patient experiences, drawing inspiration from standardized patients in medical education Barrows1993AnOO .
+We conduct systematic evaluation with senior physician to validate the effectiveness of patient agent.
+The details can be found in Appendix A.2.2.
 3.4 Benchmark Evaluation
-Given the complexity of medical decision-making, a simple binary evaluation of correct or incorrect is inadequate.
-To quantify the quality of CDM results, we’ve developed an evaluation system for each stage.
+Given the complexity of medical decision-making, we developed a comprehensive evaluation framework that goes beyond simple binary assessments.
 The standard answer for each stage comprises multiple key points.
 The evaluated agent’s score increases with the number of standard points addressed, while irrelevant content leads to score reduction.
-To implement this approach, we employ the Intersection over Union (IoU) metric.
-For instance, in the treatment stage, the agent’s score improves as it proposes more reasonable treatment measures.
+For each task, we employ specific evaluation metrics: Task 1 (Specialty Referral) uses accuracy and Intersection over Union (IoU) to compare predictions with ground truth.
+Task 2 (History-taking) calculates IoU between predicted and ground truth examination items.
+Task 3 leverages DocLens xie2024doclens for assessing image interpretation quality.
+Task 4 employs a carefully designed prompt based on National Health Commission of China guidelines to evaluate diagnostic accuracy.
+Task 5 measures IoU between predicted and ground truth examination items.
+We elaborate the evaluation metrics in Appendix A.
 4 MedChain-Agent Framework
 4.1 Multi-agent System with Feedback
-The MedChain-Agent framework introduces a novel multi-agent system designed to simulate the complex, interconnected nature of medical decision-making processes.
-This system leverages the collective intelligence of multiple specialized agents, each contributing unique expertise to the overall decision-making process.
-4.1.1 Agent Role
-Our framework incorporates three distinct agent types, each fulfilling crucial roles in the medical decision-making pipeline:
-General Agents: These specialized agents are recruited based on the specific requirements of each task.
-For instance, in Task 1 Specialty Referral, General Agents possess comprehensive knowledge of medical specialties and triage protocols, enabling them to effectively assess and route patients to appropriate departments.
-These agents engage in group discussions, simulating the collaborative nature of medical consultations.
-Summarizing Agent: In each stage of the process, a summarizing agent consolidates the discussions and analyses of the general agents.
-This agent synthesizes the collective expertise into a coherent summary and delivers the final decision for each task, mimicking the role of a senior physician or department head within a medical context.
-Feedback Agent: The feedback agent plays a crucial role in maintaining the quality and accuracy of decisions throughout the process.
-This agent evaluates the output of each task, providing constructive feedback and suggesting iterative improvements when necessary.
-By doing so, it creates a dynamic, self-correcting system that continually refines its decision-making capabilities.
-4.1.2 Decision Making
-With three types of agents, MedChain-Agent conducts a sophisticated decision-making process that mirrors real-world medical practice.
-General agents analyze patient information and engage in collaborative discussions, sharing assessments and recommendations.
-The summarizing agent then consolidates these insights, weighing different perspectives to formulate a preliminary decision.
-The feedback agent reviews this decision, evaluating its appropriateness and implications for subsequent tasks.
-If issues are identified, the feedback agent provides critiques and suggestions, initiating an iterative refinement process.
-General agents reconsider their assessments based on this feedback, while the summarizing agent refines the decision.
-This cycle continues until consensus is reached or a predefined number of iterations is completed, at which point the final decision is issued.
+The MedChain-Agent framework introduces a multi-agent system that simulates the complex, interconnected nature of medical decision-making.
+This system integrates specialized agents that contribute distinct expertise to the diagnostic and treatment process.
+We sketch the main content here and more details please refer to Appendix B.
+4.1.1 Agent Roles
+Our framework comprises three agent types:
+General Agents: Task-specific agents recruited based on domain requirements.
+For specialty referral tasks, these agents possess expertise in medical specialties and triage protocols, enabling effective patient routing.
+They engage in collaborative discussions that mirror real-world medical consultations.
+Summarizing Agent: Acts as a central coordinator, consolidating insights from general agents into coherent decisions.
+This agent synthesizes collective expertise and delivers final recommendations, similar to a senior physician’s role in clinical settings.
+Feedback Agent: Maintains decision quality through continuous evaluation and refinement.
+This agent assesses outputs, provides targeted feedback, and initiates improvements when necessary, creating a self-correcting decision process.
+4.1.2 Decision Making with Feedback
+The decision-making process begins with general agents analyzing patient data and engaging in structured discussions.
+The summarizing agent then consolidates these assessments into preliminary decisions, which undergo review by the feedback agent.
+When issues are identified, the feedback agent initiates an iterative refinement cycle, with general agents reconsidering their assessments and the summarizing agent adjusting decisions accordingly.
+This process continues until reaching consensus or completing a preset number of iterations.
+4.2 MedCase-RAG
+To enhance decision-making capabilities, we developed MedCase-RAG, a specialized Retrieval-Augmented Generation system for medical applications.
+Unlike traditional medical RAG systems, our approach employs structured case representation using feature vectors that encode 12 key clinical dimensions: Age, Sex, Chief Complaint, Symptoms, Medical History, Physical Examination, Laboratory Tests, Imaging Reports, Diagnosis, Treatment Plan, Primary Department, and Secondary Department.
+This feature vectors is extracted by text-embedding-ada-002 from openai.
+When processing new cases, MedCase-RAG performs similarity searches, identifying the three most similar cases using cosine similarity metrics.
+This approach provides agents with comprehensive reference points for decision-making.
+The MedCase-RAG also features dynamic knowledge base expansion, incorporating resolved cases as pseudo-data.
+This mechanism enables continuous learning from new clinical experiences, adapting to evolving medical knowledge and practices.
 Table 1: Evaluation of various LLM-based agent in MedChain. The best performance for each task is highlighted in bold.
@@
 Level 1 Level 2
-Single-agent GPT-4o-mini 2023 0.5449 0.2871 0.3399 0.5112 0.4406 0.3930 0.4195
-Claude-3.5-sonnet 2023 0.5681 0.3050 0.3562 0.5018 0.4599 0.4053 0.4327
-LLaVA 2024a 0.3240 0.0730 0.3182 0.5165 0.4024 0.1060 0.2901
-Qwen2 2024a 0.4975 0.2215 0.4226 0.4829 0.4818 0.2193 0.3876
-InternVL2 2024a 0.4811 0.1935 0.4645 0.4490 0.4738 0.2903 0.3920
-Multi-agent MedAgent 2023 0.3830 0.2039 0.4454 0.4456 0.4266 0.3673 0.3786
-MDAgent 2024a 0.2398 0.1343 0.4240 0.4983 0.4289 0.3620 0.3479
-MedChain-Agent 0.5873 0.3505 0.5836 0.6566 0.5218 0.4613 0.5269
-4.2 MedCase-RAG
-To enhance the decision-making capabilities of our multi-agent framework, we introduce MedCase-RAG, a novel Retrieval-Augmented Generation technique tailored for medical contexts.
-Unlike conventional medical RAG methods that rely on unstructured medical QA databases, MedCase-RAG employs a structured approach to data representation.
-We map each medical case into a 12-dimensional feature vector, encompassing crucial attributes such as Age, Sex, Patient Description, Symptom Description, Patient History, and Patient Image.
-This structured representation allows for more comprehensive and standardized case encoding, facilitating more accurate and nuanced retrieval.
-Within this feature set, we identify "Symptom Description" as the most representative feature of a patient’s condition.
-This feature undergoes quantification through a Text Embedding model, serving as the primary key for dense retrieval tasks.
-Moreover, MedCase-RAG incorporates a dynamic database expansion mechanism to continually enrich its knowledge base.
-As patients complete their medical processes and recover, their case information is reintroduced into the database as pseudo-data.
-This approach allows the system to learn from new cases and adapt to evolving medical knowledge and practices.
-When a new patient case is presented, MedCaseRAG performs a similarity search within the relevant medical department, using cosine similarity to identify the top three most similar cases, offering a more comprehensive basis for decision-making, allowing agents to consider various aspects and potential outcomes.
+Base LLM
+GPT-4o-mini openai2023gpt4v 0.5449 0.2871 0.3399 0.5112 0.4177 0.3930 0.4156
+GPT-o3 0.5495 0.2866 0.3493 - 0.4891 0.3343 0.4017
+Claude-3.5-sonnet team2023gemini 0.5681 0.3050 0.3562 0.5018 0.4207 0.4053 0.4262
+MedGemma sellergren2025medgemma 0.5063 0.1975 0.3315 0.6324 0.4668 0.3558 0.4105
+LLaVA liu2024visual 0.3240 0.0730 0.3182 0.5165 0.3667 0.1060 0.2841
+Qwen2 wang2024qwen2 0.4975 0.2215 0.4226 0.4829 0.4530 0.2193 0.3828
+InternVL2 wang2024qwen2 0.4811 0.1935 0.4645 0.4490 0.4367 0.2903 0.3859
+Baichuan dou2025baichuan 0.2959 0.0532 0.3264 - 0.4297 0.2591 0.2728
+HuaTuoGPT zhang2023huatuogpt 0.0707 0.0207 0.3465 - 0.4163 0.1780 0.2064
+MedReason wu2025medreason 0.4602 0.1258 0.3315 - 0.4741 0.2993 0.3381
+FineMedLM-o1 yu2025finemedlm 0.1502 0.0124 0.3969 - 0.2772 0.1345 0.1942
+Single-agent
+Zero-shot 0.4811 0.1935 0.3450 0.4994 0.4572 0.2817 0.3763
+Few-shot 0.5584 0.2481 0.4870 0.3758 0.4382 0.3553 0.4105
+CoT wei2022chain 0.5698 0.1484 0.3750 0.6396 0.4382 0.3421 0.4189
+RAG lewis2020retrieval 0.5927 0.2467 0.4732 0.6408 0.4167 0.4524 0.4704
+Self-consistency wang2022self 0.5143 0.2319 0.4213 0.4144 0.4198 0.3441 0.3910
+Multi-agent DyLAN liu2023dynamic 0.4415 0.1731 0.4434 0.4980 0.3863 0.2983 0.3734
+AutoGen wu2023autogen 0.5228 0.2347 0.4528 0.4559 0.4250 0.3590 0.4084
+MedAgent tang2023medagents 0.3830 0.2039 0.4454 0.4456 0.4102 0.3673 0.3759
+MDAgent kim2024adaptive 0.2398 0.1343 0.4240 0.4983 0.3973 0.3620 0.3426
+MDAgent + RAG kim2024adaptive 0.4754 0.2063 0.4412 0.5346 0.4198 0.4371 0.4190
+MedChain-Agent 0.5873 0.3505 0.5836 0.6566 0.4807 0.4613 0.5200
 5 Experiments
@@
 We split the dataset into training, validation, and testing sets with a ratio of 7:1:2.
-For frameworks incorporating RAG techniques, both training and validation sets are used to construct the case retrieval database.
 Our study evaluates both single-agent and multi-agent systems.
-For single agents, we test two closed-source models (gpt-4o-mini OpenAI (2023), and claude-3.5-sonnet Team et al.
-(2023)) and three open-source models (InternVL2-8b Chen et al.
-(2023), llava-llama-3-8b-v1_1 Liu et al.
-(2024a), and Qwen2-7B-Instruct Wang et al.
-(2024a)), with model weights obtained from official Hugging Face repositories.
-In the multi-agent evaluation, we compare MedChain-Agent against MedAgent Tang et al.
-(2023) and MDAgent Kim et al.
-(2024a).
-All multi-agent framework are based on InternVL2-8b Chen et al.
-(2023).
-The deployment was conducted using the LMDeploy framework Contributors (2023).
+For base LLM, we test two closed-source models (gpt-4o-mini openai2023gpt4v , and claude-3.5-sonnet team2023gemini ) and four open-source models (InternVL2-8b chen2023internvl , llava-llama-3-8b-v1_1 liu2024visual , HuaTuoGPT zhang2023huatuogpt , and Qwen2-7B-Instruct wang2024qwen2 ), with model weights obtained from official Hugging Face repositories.
+In the single-agent evaluation, we compare with zero-shot manner, few-shot manner, CoT wei2022chain , Self-consistency wang2022self and RAG lewis2020retrieval .
+In the multi-agent evaluation, we compare MedChain-Agent against DyLAN liu2023dynamic , AutoGen wu2023autogen , MedAgent tang2023medagents and MDAgent kim2024adaptive .
+All agent framework are based on InternVL2-8b chen2023internvl .
+The deployment was conducted using the LMDeploy framework 2023lmdeploy .
 All tests executed on NVIDIA A100 GPUs featuring 80GB of memory.
 To enhance output stability and reliability across all experiments, we consistently set the temperature parameter to 0.
+The experiments were conducted in Chinese.
+Table 2: Performance Comparison of LLM-based Agents across Other Diagnosis Benchmarks.
+Implementation of agent frameworks based on InternVL2, evaluated on multiple medical benchmarks including MedQA jin2021disease , PubMedQA jin2019pubmedqa , PathVQA he2020pathvqa , and MedBullets chen2024benchmarking .
+Framework Method MedQA PubMedQA PathVQA MedBullets Average
+Single-agent Zero-shot 0.426 0.668 0.449 0.490 0.508
+Few-shot 0.477 0.648 0.448 0.503 0.519
+CoT 0.470 0.714 0.465 0.500 0.537
+Self-consistency 0.460 0.688 0.482 0.500 0.533
+Multi-agent MedAgents 0.501 0.622 0.569 0.435 0.532
+MDAgents 0.435 0.744 0.582 0.422 0.546
+AutoGen 0.395 0.656 0.568 0.448 0.517
+DyLAN 0.414 0.610 0.540 0.448 0.503
+MedChain-Agent 0.462 0.746 0.621 0.474 0.576
+5.2 Benchmark Performance Results
+The results of our evaluation in the MedChain  are presented in Table 1. Our analysis yields two significant insights:
+(1) Sequential decision-making tasks continue to pose significant challenges, even for advanced models.
+For instance, within the single-agent frameworks, GPT-4o-mini and InternVL2 achieve average scores of 0.4156 and 0.3859, respectively.
+These results indicate that despite their sophistication, these models struggle to maintain consistent performance across the sequential stages of clinical decision-making, highlighting the inherent difficulty of these tasks.
+(2) The integration of the MedChain-Agent framework with open-source LLMs demonstrates significant superiority over proprietary models like GPT-4o-mini.
+The substantial performance gain observed with MedChain-Agent (average score of 0.5200) implies that our framework can leverage the strengths of open-source LLMs to achieve superior outcomes.
+This suggests that open-source models, when enhanced with our framework, are not only competitive but can also excel in handling intricate medical decision-making tasks.
+5.3 Evaluation in Existing Diagnosis Datasets
+Moreover, we evaluate MedChain-Agent on several well-established medical QA datasets, including MedQA jin2021disease , PubMedQA jin2019pubmedqa , PathVQA he2020pathvqa , and MedBullets chen2024benchmarking .
+As shown in Table 2, our framework demonstrates strong performance compared to both single-agent baselines and multi-agent alternatives.
+MedChain-Agent achieves the highest average score (0.576).
+This consistent superiority across various medical QA benchmarks further validates the effectiveness of our framework, even on simpler, more structured tasks that differ from our real-world clinical scenarios.
+5.4 Ablation Studies and Discussion
 Table 3: Ablation Study for Key Components for MedChain-Agent.
@@
 Level 1 Level 2
-0.5523 0.2228 0.3285 0.6369 0.4982 0.3915 0.4384
-0.5739 0.2906 0.4222 0.6377 0.5086 0.4209 0.4757
-0.5928 0.3353 0.5801 0.6488 0.5035 0.4568 0.5195
-0.5873 0.3505 0.5836 0.6566 0.5218 0.4613 0.5269
-5.2 Evaluation Metrics
-To assess model performance, we employ diverse metrics tailored to each task.
-Accuracy is used for tasks with a single correct answer, such as Specialty Referral (Level 1).
-For tasks involving multiple key points, including Specialty Referral (Level 2), History-taking, and Treatment, we apply Intersection over Union (IoU) to capture the overlap between predicted and ground truth sets.
-Specifically for the Examination task, we utilize DocLens Xie et al.
-(2024) to evaluate the quality of image interpretation reports.
-5.3 Experiment Result and Analysis
-The results of our evaluation in the MedChain  are presented in Table 2. Our analysis yields three significant insights:
-(1) Sequential decision-making tasks continue to pose significant challenges, even for advanced models.
-For instance, within the single-agent frameworks, GPT-4o-mini and InternVL2 achieve average scores of 0.4442 and 0.3920, respectively.
-These results indicate that despite their sophistication, these models struggle to maintain consistent performance across the sequential stages of clinical decision-making, highlighting the inherent difficulty of these tasks.
-(2) Multi-agent frameworks based on InternVL2, such as MedAgent and MDAgent, exhibit inferior performance compared to their single-agent counterparts.
-This degradation suggests that traditional multi-agent approaches may exacerbate error propagation in sequential decision-making processes, leading to reduced overall performance.
-In contrast, our proposed MedChain-Agent significantly outperforms these multi-agent methods, achieving an average score of 0.5269.
-This improvement demonstrates that MedChain-Agent effectively mitigates error propagation, enhancing decision quality and reliability in clinical settings.
-(3) The integration of the MedChain-Agent framework with open-source LLMs demonstrates significant superiority over proprietary models like GPT-4o-mini.
-The substantial performance gain observed with MedChain-Agent (average score of 0.5269) implies that our framework can leverage the strengths of open-source LLMs to achieve superior outcomes.
-This suggests that open-source models, when enhanced with our framework, are not only competitive but can also excel in handling intricate medical decision-making tasks.
-5.4 Ablation Study
+0.5523 0.2228 0.3285 0.6369 0.4724 0.3915 0.4341
+0.5739 0.2906 0.4222 0.6377 0.4299 0.4209 0.4692
+0.5928 0.3353 0.5801 0.6488 0.4699 0.4568 0.5140
+0.5873 0.3505 0.5836 0.6566 0.4804 0.4613 0.5200
+Table 4: Generalizability of MedChain-Agent with Various Base LLM.
+We apply the MedChain-Agent framework to different base LLMs to validate its generalizability.
+Performance changes are highlighted with light green for improvements and light red for decreases.
+Results demonstrate that the MedChain-Agent framework consistently brings performance gains across different base LLMs.
+Base LLM Specialty Referral History-taking Examination Diagnosis Treatment Average
+Level 1 Level 2
+GPT-4o-mini 0.6065 (+0.0616) 0.2302 (-0.0569) 0.6205 (+0.2806) 0.6279 (+0.1167) 0.4975 (+0.0798) 0.4329 (+0.0399) 0.5026 (+0.0870)
+HuaTuoGPT 0.2554 (+0.1847) 0.0097 (-0.0110) 0.3722 (+0.0257) - 0.4159 (-0.0004) 0.1450 (-0.0330) 0.2396 (+0.033)
+Qwen2 0.5818 (+0.0843) 0.2781 (+0.0566) 0.5962 (+0.1736) 0.6534 (+0.1705) 0.4628 (+0.0098) 0.4315 (+0.2122) 0.5006 (+0.1178)
+InternVL2 0.5873 (+0.1062) 0.3505 (+0.1570) 0.5836 (+0.1191) 0.6566 (+0.2076) 0.4806 (+0.0439) 0.4613 (+0.1710) 0.5200 (+0.1341)
+Table 5: The ablation study for three key characteristics in MedChain.
+This table presents the impact of personalization, interactivity, and sequentiality on Diagnosis and Treatment tasks.
+The arrows $\uparrow\downarrow$ next to settings indicate expected performance change direction.
+The arrows next to results show actual changes (highlighted in gray when matching expectations).
+Setting Model Diagnosis Treatment
+Full MedAgent 0.4106 0.3673
+MDAgent 0.3959 0.3620
+gpt-4o-mini 0.4157 0.3930
+InternVL2 0.4378 0.4472
+MedChain-Agent 0.4802 0.4613
+a) w/o Person. $\downarrow$ MedAgent 0.3075 $\downarrow$ 0.3754 $\uparrow$
 […]
+[…]
+[…]
+MedChain-Agent 0.4159 $\downarrow$ 0.4310 $\downarrow$
 Setting Model Diagnosis Treatment
-Full MedAgent 0.4266 0.3673
-MDAgent 0.4289 0.3620
-gpt-4o-mini 0.4406 0.3930
-InternVL2 0.4759 0.4472
-w/o Person. $\downarrow$ MedAgent 0.3274 $\downarrow$ 0.3754 $\uparrow$
 […]
@@
 […]
+MedChain-Agent 0.4807 $\uparrow$ 0.4743 $\uparrow$
+c) w/o Inter. $\uparrow$ MedAgent 0.3129 $\downarrow$ 0.4109 $\uparrow$
+MDAgent 0.3998 $\uparrow$ 0.3627 $\uparrow$
 […]
-[…]
-[…]
-w/o Inter. $\uparrow$ MedAgent 0.3315 $\downarrow$ 0.4109 $\uparrow$
-MDAgent 0.4295 $\uparrow$ 0.3627 $\uparrow$
-[…]
-[…]
-Three key characteristics for MedChain: To validate the effectiveness of personalization, interactivity, and sequentiality within our benchmark, we conduct ablation study as in Table 4.
+InternVL2 0.4550 $\uparrow$ 0.3173 $\downarrow$
+MedChain-Agent 0.4634 $\downarrow$ 0.5207 $\uparrow$
+(1) Ablation Study for Key components in MedChain-Agent: To assess our framework’s components, we conduct ablation studies as shown in Table 3.
+Both the Feedback mechanism and MedCase-RAG module demonstrate significant individual contributions to overall performance.
+The Feedback mechanism alone improves the average score from 0.4341 to 0.4692, with notable gains in History-taking and Level 2 Specialty referral.
+MedCase-RAG shows stronger individual impact, boosting the average score to 0.5140, with substantial improvements in History-taking and Specialty referral tasks.
+While MedCase-RAG excels in diagnostic phases, the Feedback mechanism appears more beneficial for Treatment tasks.
+When combined, these components show synergistic effects, achieving the highest average performance (0.5200) and optimal scores across five of six evaluated tasks, validating their complementary roles in enhancing clinical reasoning capabilities.
+(2) Ablation Study for Three Key Characteristics in MedChain: To validate the effectiveness of personalization, interactivity, and sequentiality within our benchmark, we conduct an ablation study as shown in Table 5.
 We systematically remove each characteristic and observe its impact on model performance across Diagnosis and Treatment tasks.
-1) Removing patient-specific information (w/o Person.) consistently degrades performance across all models in diagnosis tasks (with drops ranging from 3.0% to 13.4%).
-This demonstrates that personalized information is crucial for accurate clinical decision-making.
-2) Interestingly, when removing the sequential dependency between stages (w/o Seq.), most models show improved performance.
-This improvement actually validates the effectiveness of sequentiality in our benchmark, as it indicates that sequential decision-making poses greater challenges that better reflect real-world clinical scenarios.
-3) Similarly, the improved performance observed after removing interactive information gathering (w/o Inter.) confirms the effectiveness of interactivity in our benchmark design.
+’w/o Person’ means all detailed patient profiles are omitted from the input, resulting in a lack of personalized information that makes correct diagnosis more challenging for the model.
+’w/o Seq’ means we use the ground truth from the previous stage as input to the next stage, rather than using the model’s previous output, which simplifies the benchmark.
+’w/o Inter’ means we directly provide all patient examination results as input without requiring the agent to autonomously inquire about the patient’s condition, also simplifying the benchmark.
+Removing patient-specific information (w/o Person.) consistently degrades performance across all models in diagnosis tasks (with drops ranging from 2.51% to 10.31%), demonstrating that personalized information is crucial for accurate clinical decision-making.
+Interestingly, when removing the sequential dependency between stages (w/o Seq.), most models show improved performance, indicating that sequential decision-making poses greater challenges that better reflect real-world clinical scenarios.
+Similarly, the improved performance observed after removing interactive information gathering (w/o Inter.) confirms the effectiveness of interactivity in our benchmark design.
 These results collectively suggest that while both sequentiality and interactivity make the benchmark more challenging, they are essential components that better simulate the complexity of real-world clinical decision-making processes.
-Key components for MedChain-Agent: To assess the impact of individual components in our framework, we conducted ablation experiments, with results shown in Table 3.
-These findings highlight the crucial roles of both the Feedback mechanism and the MedCase-RAG module in enhancing overall performance.
-Each component independently contributes to significant improvements across various metrics.
-Notably, the MedCase-RAG module alone boosts Specialty Referral (Level 1) to 0.5928 and yields substantial gains in History-taking and Examination.
-When combined, Feedback and MedCase-RAG demonstrate a synergistic effect, achieving an impressive average score of 0.5269.
-This underscores the complementary nature of these components in optimizing our framework’s capabilities.
+(3) Generalizability for MedChain-Agent: To evaluate the generalizability of our framework, we apply MedChain-Agent to various base LLMs as shown in Table 4.
+The results demonstrate substantial performance improvements across most models and tasks.
+InternVL2 shows the most significant enhancement with a 13.41% average improvement, followed by Qwen2 (11.78%) and GPT-4o-mini (8.70%).
+Even with HuaTuoGPT, which has relatively lower baseline performance, our framework still achieves a 3.33% improvement.
+Notably, the improvements are particularly pronounced in complex tasks such as History-taking and Examination, where gains of up to 28.06% are observed.
+These consistent enhancements across diverse models and tasks strongly validate the framework’s robust generalizability and effectiveness in medical decision-making scenarios.
 6 Conclusion
-In this paper, we introduced MedChain, a novel benchmark for evaluating large language models in clinical decision-making, and MedChain-Agent, an innovative multi-agent framework designed to address the complexities of real-world medical scenarios.
-Our work bridges a critical gap between current AI capabilities and the nuanced realities of clinical practice by incorporating personalization, sequentiality, and interactivity into the evaluation process.
-MedChain-Agent’s sophisticated multi-agent system, enhanced by a feedback mechanism and the MedCase-RAG module, demonstrates unprecedented performance in sequential clinical decision-making tasks, even outperforming proprietary models when combined with open-source LLMs.
-By providing a more realistic and challenging evaluation framework, this research sets a new standard for evaluating and developing medical AI systems, paving the way for its responsible integration into clinical practice.
-Acknowledgement
-We are particularly indebted to the administrators of the iiyi website for their generosity in allowing us to utilize their data for our research purposes.
-We would like to acknowledge the assistance provided by Claude-3.5 in proofreading our manuscript for grammatical accuracy and in facilitating the creation of LaTeX tables.
-Limitations
+In this paper, we introduced MedChain, a novel benchmark for evaluating LLM-based agents in clinical decision-making that authentically reflects real-world medical practice through three essential characteristics: personalization, interactivity, and sequentiality.
+Our comprehensive dataset encompasses 12,163 diverse clinical cases across 19 medical specialties, including 7,338 medical images with corresponding reports, providing a robust foundation for evaluating AI systems in complex healthcare scenarios.
+To address the challenges presented by this benchmark, we also introduced MedChain-Agent, a multi-agent framework enhanced by feedback mechanisms and MedCase-RAG, which demonstrates superior performance across various clinical tasks.
+This work establishes new benchmarks for evaluating medical AI systems and provides practical solutions for enhancing their clinical decision-making capabilities.
+As AI continues to evolve in healthcare applications, frameworks that can navigate the complexities of real-world clinical scenarios will be increasingly valuable for improving patient care and supporting healthcare professionals.
+7 Limitations
 This paper has two primary limitations that offer avenues for future research:
-  • Data Source Diversity: The MedChain benchmark is constructed from 12,163 electronic health records obtained from the Chinese medical website “iiYi." Although this dataset is extensive and covers 19 medical specialties and 156 sub-categories, it is derived from a single source.
-In our future work, we will incorporate additional data sources from different regions or healthcare systems to further enrich the dataset, providing a broader spectrum of clinical scenarios and enhancing the generalizability of the benchmark.
-  • Patient Interaction Simulation: In our interactive environment, the patient responses during the history-taking stage are generated by the Gemma 2 language model.
+1) Data Source Diversity: The MedChain benchmark is constructed from 12,163 electronic health records obtained from the Chinese medical website “iiYi." Although this dataset is extensive and covers 19 medical specialties and 156 sub-categories, it is derived from a single source.
+Additionally, there exists a notable imbalance in data distribution across different medical specialties, with some departments having significantly more cases than others.
+This imbalance may introduce bias in model evaluation and limit performance in underrepresented specialties.
+In our future work, we will incorporate additional data sources from different regions or healthcare systems to further enrich the dataset and address the specialty imbalance.
+2) Patient Interaction Simulation: In our interactive environment, the patient responses during the history-taking stage are generated by the Gemma 2 language model.
 While this approach ensures consistency and control in evaluating the LLM-based agent, the real patient interactions can be more varied and complex.
@@
 Appendix C lists the additional details for experiment.
-Appendix A Benchmark Construction
+Appendix D discuss the difference between MedChain and several similar related works.
+Appendix A Benchmark Construction and Evaluation
 A.1 Dataset Standardization
@@
 This standardization process ensures that the dataset is consistent, facilitating accurate and comparable evaluations of LLM performance.
-Figure 5 shows the statistics of different departments and Figure 4 demonstrate a case after standardization.
-Figure 4: Case presentation "77_Ovarian Carcinoid with Mature Cystic Teratoma: A Case Report."
+Figure 4 and Figure 5 demonstrate a case after standardization in English and Chinese.
+Figure 6 shows the statistics of different departments.
+Figure 4: Case Report "77_Ovarian Carcinoid with Mature Cystic Teratoma: A Case Report."
+Figure 5: Case Report Chinese Version with Corresponding Medical Imaging. "77_Ovarian Carcinoid with Mature Cystic Teratoma: A Case Report." (a), (b) and (c) Medical Imaging. (d) Chinese version.
 Figure 6: Multi-departmental Distribution
@@
 This includes the patient’s chief complaint, medical history, and examination results.
-  • Extraction Agent: This agent extracts examination items that the Doctor Agent inquires about or suggests from the dialogue between the Doctor and Patient Agents.
-These extracted items form the output for comparison with the ground truth examination set.
-The evaluation metric is the IoU between the examination items extracted by the Extraction Agent (based on the Doctor Agent’s inquiries and suggestions) and the ground truth set.
-This multi-agent design simulates a realistic clinical history-taking process, allowing the LLM to demonstrate its ability to ask relevant follow-up questions, interpret patient responses, and determine appropriate examinations.
+Prompt for Patient: You are to role-play as a Standardized Patient.
+Here is your case information: {patient information}.
+You need to answer the doctor’s questions directly based on the case information (do not fabricate doctor-patient dialogues).
+Note that unless the doctor explicitly asks about physical examination and auxiliary examination findings, please do not proactively mention or inquire about physical examination and auxiliary examination related content.
+If the doctor asks about content that does not exist in the case, please indicate that you don’t know and avoid fabricating information.
+At all times, remember that you are only playing the role of a standardized patient.
+Prompt for Doctor: You are a doctor.
+The patient’s chief complaint is as follows: {chief complaints}.
+You need to gather more information through conversation with the patient.
+Physical examination includes: general examination (including height, weight, temperature, blood pressure, pulse, etc.), head, eyes, ears, nose, and throat examination, neck examination (including thyroid, cervical lymph nodes), chest examination (including lungs, heart), abdominal examination, spine and limb examination, skin examination, neurological examination, and genitourinary system examination.
+Auxiliary examinations include: X-ray, MRI, CT, ultrasound, nuclear medicine imaging, hematological tests, urine tests, stool tests, endoscopic examination, and pathological examination.
+After obtaining certain information (such as past history and present illness history), please select the physical examinations and auxiliary examinations to inquire about based on the patient’s condition.
+Ask about only one or two items per conversation round.
+Please inquire about as many examination items as possible (ask at least one item each for physical examination and auxiliary examination) until you can determine the condition.
+At the end of the conversation, please say "Wishing you a speedy recovery."
+Evaluation metrics is the IoU between the predicted examination items and the ground truth set.
+This patient-agent design simulates a realistic clinical history-taking process, allowing the doctor-LLM to demonstrate its ability to ask relevant follow-up questions, interpret patient responses, and determine appropriate examinations.
 The use of a local large model as the Patient Agent ensures consistency in evaluations and improves the reproducibility of the benchmark.
+Systematic Evaluation: Furthermore, we conducted a systematic evaluation with a senior physician (10+ years clinical experience) across 10 simulated cases to validate the patient agent’s effectiveness.
+The evaluation focused on three key dimensions: medical history accuracy (match rate between simulated and original EHR data), symptom consistency (temporal coherence of symptom progression), and treatment response fidelity (accuracy of responses to medications and interventions).
+Each dimension was rated on a 5-point Likert scale (1=poor, 5=excellent).
+Results demonstrated strong performance across all metrics, with mean scores of 4.0 (variance=1.0) for match rate, 3.9 (variance=1.49) for temporal coherence, and 3.7 (variance=0.81) for treatment response accuracy.
 A.2.3 Task 3: Examination
@@
 This task tests the LLM’s capability to interpret visual medical data and articulate findings in a clear, professional manner.
-Given the complexity of medical reports, conventional NLP metrics such as BLEU Papineni et al.
-(2002) and BERTScore Zhang et al.
-(2019) are inadequate for evaluating this task, as they do not capture whether the generated report aligns with the ground truth in terms of medical claims.
-Instead, we adopt the Claim Recall metric proposed in DocLens Xie et al.
-(2024) to evaluate the completeness of the generated report.
+Evaluation metrics: Given the complexity of medical reports, conventional NLP metrics such as BLEU [37] and BERTScore [69] are inadequate for evaluating this task, as they do not capture whether the generated report aligns with the ground truth in terms of medical claims.
+Instead, we adopt the Claim Recall metric proposed in DocLens [63] to evaluate the completeness of the generated report.
 First, GPT-4o-mini is used to extract a list of claims from the original ground truth report.
@@
 The prompts used for this evaluation can be found in Figure 7.
+Figure 7: The prompt for Evaluation in Task3 Examination.
 A.2.4 Task 4: Diagnosis
@@
 The output is a diagnosis in free-text format.
-We use a five-level grading system to evaluate the accuracy of the diagnosis, ranging from completely inaccurate to completely accurate.
-This graded system allows for a more nuanced evaluation of the LLM’s diagnostic capabilities, as diagnoses can be partially correct or expressed in different ways.
+Evaluation metrics: We use a five-level grading system to evaluate the accuracy of the diagnosis, ranging from 1 to 5.
+The prompt we used in LLM-based evaluation is followed the instruction of People’s Medical Publishing House  [^4] Diagnostics.
+9th Ed.
+Beijing: People’s Medical Publishing House; 2018.
 The system accounts for the complex nature of medical diagnoses, where multiple factors may contribute to the patient’s condition.
-The five-level grading system is listed as following:
-  1. 1.
-    Completely inaccurate (diagnosis completely inconsistent with the actual diagnosis)
-  2. 2.
-    Mostly inaccurate (most of the diagnosis inconsistent with the actual diagnosis)
-  3. 3.
-    Partially accurate (part of the diagnosis consistent with the actual diagnosis, but overall large discrepancy)
-  4. 4.
-    Mostly accurate (most of the diagnosis consistent with the actual diagnosis)
-  5. 5.
-    Completely accurate (diagnosis fully consistent with the actual diagnosis)
+The prompt is shown in Figure 8.
+Figure 8: The prompt for Evaluation in Task4 Diagnosis.
 A.2.5 Task 5: Treatment
@@
 The output is a set of proposed treatment items.
-The evaluation metric is the IoU between the proposed treatments and the ground truth treatment set.
+This task tests the LLM’s ability to synthesize the gathered information and apply medical knowledge to develop an appropriate treatment plan, considering factors such as drug interactions, treatment protocols, and patient-specific considerations.
+Evaluation metrics is the IoU between the proposed treatments and the ground truth treatment set.
 Since multiple treatment options may be appropriate, IoU allows for partial credit when the LLM suggests a subset of the recommended treatments or proposes additional reasonable treatments that are not part of the ground truth.
-This task tests the LLM’s ability to synthesize the gathered information and apply medical knowledge to develop an appropriate treatment plan, considering factors such as drug interactions, treatment protocols, and patient-specific considerations.
 Figure 9: Comparison between our benchmark method and previous benchmarks.
 In Figure (a), the benchmark methods used in previous work are shown, which are based on specific domain problems and cannot achieve a complete end-to-end medical task workflow.
-In Figure (b), our proposed benchmark is presented, where we effectively accomplish the full workflow of medical tasks through our proposed MCase-RAG method and the ACFM mechanism.
-Figure 7: The prompt for Extract Reference Claim and Entailment Evaluation.
-A.3 Examination Items
+In Figure (b), our proposed benchmark is presented, where we effectively accomplish the full workflow of medical tasks through our proposed MedCase-RAG method and the ACFM mechanism.
+A.3 Examination Items in Task 2
 Physical Examination: General examination (including height, weight, temperature, blood pressure, pulse, etc.), head, eyes, ears, nose and throat examination, neck examination (including thyroid, cervical lymph nodes), chest examination (including lungs, heart), abdominal examination, spine and limb examination, skin examination, neurological examination, urogenital system examination.
 Auxiliary Examinations: X-ray, MRI, CT, ultrasound, nuclear medicine imaging, blood tests, urine tests, stool tests, endoscopy, pathological examination.
-A.4 Treatment Items
+A.4 Treatment Items in Task 5
 Surgery, interventional therapy, medication, chemotherapy, antibiotic therapy, radiation therapy, physical therapy, immunotherapy, psychological therapy, traditional Chinese medicine, gene therapy.
@@
 When a problem arises in the first task, subsequent tasks will analyze and make decisions based on erroneous results, potentially leading to severe impacts on the entire process.
-To address this issue, we have introduced a feedback mechanism within the full-process framework, tightly connecting the current task with all subsequent tasks (as shown in Figure 6).
+To address this issue, we have introduced a feedback mechanism within the full-process framework, tightly connecting the current task with all subsequent tasks.
 After the current task is completed, the output results are evaluated by the physicians of the subsequent tasks (Feedback Judgement Group).
@@
 B.3 RAG Details
-We performed data restructuring on the original dataset, where each case is mapped into a 12-dimensional feature vector using attributes like Age, Sex, Patient Description, Symptom Description, Patient History, and Patient Image.
+We performed data restructuring on the original dataset, where each case is mapped into feature vectors from 12 dimensions.
 Among these, "Symptom Description" is identified as the most representative feature of the current patient and is processed through a Text Embedding model for quantification, which is stored in the database for subsequent dense retrieval tasks.
 Previous Medical-RAG methods relied on medical question-answering (QA) databases and predominantly used chunked indexing for retrieval.
-Our approach differs in several ways (as shown in Figure 5).
+Our approach differs in several ways.
 Firstly, the content of our foundational medical database is distinct; while most methods have built their databases using medical QA data, ours utilizes a tree structure.
-We initially categorize patient information according to primary medical departments (in the experimental section, we categorized into 19 primary departments) and then extract and map patient information into a two-dimensional feature representation, as illustrated in Figure 4.
+We initially categorize patient information according to primary medical departments (in the experimental section, we categorized into 19 primary departments) and then extract and map patient information into a two-dimensional feature representation.
 This storage design allows for better retention of the patient’s crucial clinical information to assist in decision-making.
@@
 Figure 11: The prompt for Task 1 in the MedChain-Agent .
+Figure 12: The prompt for Task 1 in the MedChain-Agent in Chinese.
 Figure 13: The prompt for Task 2 in the MedChain-Agent .
+Figure 14: The prompt for Task 2 in the MedChain-Agent in Chinese.
 Figure 15: The prompt for Task 3 in the MedChain-Agent .
+Figure 16: The prompt for Task 3 in the MedChain-Agent in Chinese.
 Figure 17: The prompt for Task 4 and Task5 in the MedChain-Agent .
+Figure 18: The prompt for Task 4 and Task5 in the MedChain-Agent in Chinese.
+Figure 19: The case study among different multi-agent for task1.
+Figure 20: The case study among different multi-agent for task 1in Chinese.
+Figure 21: The case study among different multi-agent for task2.
+Figure 22: The case study among different multi-agent for task2 in Chinese.
+Figure 23: The case study among different multi-agent for task3.
+Figure 24: The case study among different multi-agent for task3 in Chinese.
+Figure 25: The case study among different multi-agent for task4.
+Figure 26: The case study among different multi-agent for task4 in Chinese.
+Figure 27: The case study among different multi-agent for task5.
+Figure 28: The case study among different multi-agent for task5 in Chinese.
+Figure 29: The prompt of task 1 to 5 for other baseline.
+Table 6: Comparison of Benchmarks for LLM-based Agent.
+The MedChaincomposed of 12, 163 case with five sequential stages of the clinical workflow.
+It provide interaction environment for LLM-based agent evaluation.
+Benchmark #Num Modality Personal Information Sequential Evaluation Interaction
+MedQA [70] 1,273 Text
+PubMedQA [20] 500 Text
+MedMCQA [36] 193,155 Text
+MMedBench [39] 53,566 Text
+MedBench [29] 300,901 Text
+Asclepius [54] 3,232 Text, Imaging
+GMAI-MMBench [5] 26,000 Text, Imaging
+MedChain 12,163 Text, Imaging
 Appendix C Supplementary for Experiment
@@
 We extracted 12-dimensional features from each training set case and stored them in our Medical Dataset for subsequent retrieval tasks.
+Appendix D Related Work Discussion
+We compare MedChain with most of related work in Table 6 and Figure 9. Especially, we analyze MedChain with AI Hospital and CoD as following:
+AI Hospital [11]: While both AI Hospital and our work aim to evaluate LLMs in clinical scenarios through multi-agent interactions, there are several key distinctions.
+Firstly, MedChain significantly expands the scale and diversity of medical cases, comprising 12,163 cases across 19 specialties and 156 subspecialties, including 7,338 medical images.
+This represents a substantial advancement over AI Hospital’s dataset of 506 cases with limited specialty coverage, enabling more comprehensive evaluation of LLMs’ medical capabilities across diverse clinical scenarios.
+Secondly, MedChain introduces a more sophisticated multi-agent collaboration mechanism.
+While AI Hospital employs a basic agent interaction model, our framework incorporates a feedback-driven multi-agent system enhanced by the MedCase-RAG module.
+This module enables dynamic knowledge base expansion and supports case-based reasoning through structured feature representation.
+CoD [4] introduces an interactive approach to enhance the interpretability of medical diagnosis.
+While both CoD and our work incorporate interactive components, they differ substantially in both design objectives and implementation mechanisms.
+CoD primarily focuses on improving the interpretability of the diagnostic phase through confidence-driven interactions that demonstrate the reasoning process.
+In contrast, MedChain takes a more comprehensive approach by simulating the complete clinical workflow through a multi-agent collaborative framework.
+Our framework encompasses five sequential stages from triage to treatment, with each stage’s decisions being evaluated and guided by subsequent stages through a novel cross-stage feedback mechanism.
+This design enables MedChain to capture the interdependent nature of clinical decision-making, where decisions at each stage influence and are influenced by other stages in the workflow.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: See abstract and Section 1.
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  5. 2.
+    Limitations
+  6. Question: Does the paper discuss the limitations of the work performed by the authors?
+  7. Answer: [Yes]
+  8. Justification: See section 7.
+  9. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  10. 3.
+    Theory assumptions and proofs
+  11. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  12. Answer: [N/A]
+  13. Justification: The paper does not include theoretical results.
+  14. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  15. 4.
+    Experimental result reproducibility
+  16.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  17. Answer: [Yes]
+  18. Justification: See section 5 and Appendix A, B, and C.
+  19. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  20. 5.
+    Open access to data and code
+  21. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  22. Answer: [Yes]
+  23. Justification: Dateset is uploaded in https://huggingface.co/datasets/ljwztc/MedChain. Code will be released in https://github.com/ljwztc/MedChain.
+  24. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  25. 6.
+    Experimental setting/details
+  26. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  27. Answer: [Yes]
+  28. Justification: See section 5 and Appendix A, B, and C.
+  29. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  30. 7.
+    Experiment statistical significance
+  31. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  32. Answer: [N/A]
+  33. Justification: All model weights are obtained from their official repositories on Hugging Face to ensure consistency and reliability.
+  34. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  35. 8.
+    Experiments compute resources
+  36. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  37. Answer: [Yes]
+  38. Justification: See section 5.1.
+  39. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  40. 9.
+    Code of ethics
+  41. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  42. Answer: [Yes]
+  43. Justification: The research fully conforms to the NeurIPS Code of Ethics, adhering to all ethical guidelines without deviation.
+  44. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  45. 10.
+    Broader impacts
+  46. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  47. Answer: [Yes]
+  48. Justification: See in section 6.
+  49. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  50. 11.
+    Safeguards
+  51.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  52. Answer: [N/A]
+  53. Justification: Not applicable.
+  54. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  55. 12.
+    Licenses for existing assets
+  56.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  57. Answer: [Yes]
+  58. Justification: We obtain the permission from the website.
+  59. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  60. 13.
+    New assets
+  61. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  62. Answer: [Yes]
+  63. Justification: See the abstract for a link to the dataset, website, and the code, which include details about our new benchmark
+  64. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  65. 14.
+    Crowdsourcing and research with human subjects
+  66.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  67. Answer: [No]
+  68. Justification: Not applicable.
+  69. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  70. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  71.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  72. Answer: [No]
+  73. Justification: Not applicable.
+  74. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  75. 16.
+    Declaration of LLM usage
+  76.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  77. Answer: [Yes]
+  78. Justification: We test the performance of LLM in our dataset.
+  79. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.